### v3

# Effective Embodied Representation

Abstract
In this paper, we introduce SPA, a novel representation learning framework that emphasizes the importance of 3D spatial awareness in embodied AI. Our approach leverages differentiable neural rendering on multi-view images to endow a vanilla Vision Transformer (ViT) with intrinsic spatial understanding. We present the most comprehensive evaluation of embodied representation learning to date, covering 268 tasks across 8 simulators with diverse policies in both single-task and language-conditioned multi-task scenarios. The results are compelling: SPA consistently outperforms more than 10 state-of-the-art representation methods, including those specifically designed for embodied AI, vision-centric tasks, and multi-modal applications, while using less training data. Furthermore, we conduct a series of real-world experiments to confirm its effectiveness in practical scenarios. These results highlight the critical role of 3D spatial awareness for embodied representation learning. Our strongest model takes more than 6000 GPU hours to train and we are committed to open-sourcing all code and model weights to foster future research in embodied representation learning. Project Page: https://haoyizhu.github.io/spa/.
Figure 1: Performance comparison across representations. Above: (a) Mean rank and (b) mean success rate on benchmarks. Lines represent the performance of SPA, best, and second best performance on each benchmark. Bottom: Rank distributions for 268 individual tasks, showing proportions from rank 1 to rank $\geq$ 4 counterclockwise. Our model demonstrates superior overall performance.
1 Introduction
Vision systems have made remarkable progress in understanding 2D images (He et al., 2020; Chen et al., 2020a; He et al., 2022; Feichtenhofer et al., 2022; Tong et al., 2022; Yang et al., 2023; Oquab et al., 2023; Radford et al., 2021; Fang et al., 2023b; Chen et al., 2024b). However, achieving true visual intelligence necessitates a comprehensive understanding of the 3D world. This is crucial for embodied AI, where agents must perceive, reason, and interact with complex 3D environments.
Existing visual representation learning methods for embodied AI (Nair et al., 2022; Radosavovic et al., 2023; Majumdar et al., 2023; Karamcheti et al., 2023; Shang et al., 2024; Yang et al., 2024b) largely rely on paradigms from 2D vision, predominantly employing contrastive-based or masked autoencoder (MAE)-based approaches. However, they often struggle to fully capture the spatial relationships and 3D structures inherent in the physical world. This limitation arises from their primary emphasis on 2D semantic understanding, which, though valuable, is still insufficient for the sophisticated spatial reasoning required in embodied AI tasks, where agents need to navigate environments, manipulate objects, and make decisions using their 3D spatial awareness.
In this paper, we introduce SPA, a general 3D spatial-aware representation learning framework for embodied AI. SPA leverages neural rendering (Mildenhall et al., 2021) as the pre-training pre-text task on multi-view images. Unlike explicit 3D representations like point clouds or meshes—which prior work (Wang et al., 2024b; a; Ze et al., 2024; Zhu et al., 2024) has shown to outperform pure 2D inputs in robot learning—multi-view images are easier to process and more readily available, making them ideal for large-scale training, such as from internet videos. Specifically, given a vanilla 2D image backbone, e.g. a Vision Transformer (ViT) (Dosovitskiy et al., 2021), we first extract multi-view feature maps from the input images. Using known camera poses, we then construct a feature volume from these feature maps and sample rays to apply differentiable neural rendering. This process generates multi-view RGB-D images and semantic maps for supervision without labels, enabling the pre-training of a 2D image backbone to enhance 3D spatial awareness.
To thoroughly validate our assumption and method, we collect 268 embodied tasks across 8 simulators using various policy methods. To our knowledge, this represents the largest scale of embodied evaluation to date. Previous work, such as R3M (Nair et al., 2022) and VC-1 (Majumdar et al., 2023), evaluated fewer than 20 tasks, potentially leading to incomplete or biased conclusions. Our evaluation spans both single-task and language-conditioned multi-task learning. We compare over 10 state-of-the-art representation learning methods, categorized as embodied-specific (Nair et al., 2022; Majumdar et al., 2023; Radosavovic et al., 2023), vision-centric (Oquab et al., 2023; Chen et al., 2021; He et al., 2022), and multi-modal (Radford et al., 2021; Fang et al., 2023b; Chen et al., 2024b). Our method consistently outperforms others, underscoring the importance of 3D spatial awareness for embodied AI. Notably, multi-modal models like CLIP (Radford et al., 2021), consistently perform poorly. This holds even the vision-language model scales the ViT to 6B parameters (Chen et al., 2024b). Through a camera pose estimation task and feature map visualization, we demonstrate that SPA has learned superior 3D spatial understanding. Further, we find that 3D awareness shows a positive correlation with embodied performance. Finally, we conduct several real-world tasks, where SPA also demonstrates superior performance. Our contribution can be summarized as follows.
  • We propose a significant spatial hypothesis: 3D spatial awareness is crucial for embodied representation learning. Our experiments provide clear evidence for the hypothesis.
  • We introduce SPA, a novel paradigm for representation learning in embodied AI. It enhances a vanilla Vision Transformer (ViT) with 3D awareness using differentiable neural rendering as the pre-text task on multi-view images.
  • We conduct the largest evaluation benchmark for embodied representation learning, significantly larger than previous studies. It involves 268 tasks, 8 simulators, and over 10 SOTA methods with diverse downstream policies and task settings.
  • Through extensive experiments in both simulators and real-world settings, SPA outperforms more than 10 SOTA representation learning methods, demonstrating its effectiveness.
2 Methodology
In this section, we first describe our process for handling multi-view image inputs and feature extraction in Sec. 2.1. Subsequently, we construct an explicit feature volume from these multi-view features, detailed in Sec. 2.2. Finally, we explain the image rendering from the feature volume and loss functions for network optimization in Sec. 2.3 and Sec. 2.4. Our pipeline is visualized in Fig. 2.
2.1 Input Process and Feature Extraction
Given a set of multi-view images $\mathbf{I}=\{I_{1},I_{2},\ldots,I_{N}\}$, where each $I_{i}\in\mathbb{R}^{3\times H\times W}$ and $N\in\mathbb{Z}^{+}$, we utilize a 2D image backbone $F$, such as a ViT. The images are processed separately through $F$, yielding latent features $\mathbf{L}=\{l_{1},l_{2},\ldots,l_{N}\}$, where each $l_{i}=F(I_{i})\in\mathbb{R}^{L\times C}$. Following MAE, we apply random masking to input images to enhance robustness, but without a ViT decoder and MAE’s pixel reconstruction objective. For each $l_{i}$, masked positions are filled with a mask token, and we concatenate the global class token with other patch tokens as read-out tokens similar to DPT (Ranftl et al., 2020). We then unpatchify them to obtain a latent feature map of size $\frac{H}{P}\times\frac{W}{P}$, where $P$ is the ViT patch size. Finally, two simple upsampling layers transform this into a feature map $M_{i}$ matching the input resolution. Each upsampling layer includes a convolution, a GELU (Hendrycks & Gimpel, 2016) activation, and a pixel shuffle layer (Shi et al., 2016) with an upscale factor of $\sqrt{P}$.
2.2 Dynamic Volume Construction
To enable multi-view interaction, we construct a 3D feature volume from multi-view feature maps, $\mathbf{M}$. Unlike the bird’s-eye view (BEV) construction in autonomous driving (Li et al., 2022), which usually relies on a fixed scene range around ego vehicle , our method dynamically adjusts the scene range based on the spatial extents of the environment to accommodate varying datasets. Specifically, the scene’s bounds are first estimated using available depth data, sparse points, or pre-defined rules. We then partition the scene into a volume of size $X\times Y\times Z$, with voxel size dynamically adjusted to capture either fine object details or larger environments. Voxel features, $\tilde{\mathcal{V}}$, are initialized with learnable positional embeddings. Each voxel is projected onto the multi-view feature maps using the known transformation matrix $\mathbf{T}$. Deformable attention (Zhu et al., 2021) is then applied, where the multi-view features act as keys and values, and the voxel features as queries. Finally, a 3D convolution refines the output volume features to obtain $\mathcal{V}$. The process can be formulated as:
$\mathcal{V}=\operatorname{Conv3D}(\operatorname{DeformAttn}(\tilde{\mathcal{V}},\mathbf{M},\mathbf{T})).$ (1)
Figure 2: Pipeline Overview. Given multi-view images, we randomly mask patches and input the remaining into a Vision Transformer. The upsampled latent features generate multi-view feature maps, from which we construct a feature volume to derive SDF values, SH coefficients, and semantic features. We then render depth, RGB, and semantic maps for loss computation.
2.3 Differentiable Volumetric Rendering
After constructing the feature volume, we employ differentiable neural rendering (Mildenhall et al., 2021) to connect 2D and 3D domains. For better geometry representation, we utilize the implicit signed distance function (SDF) field modeling as in NeuS (Wang et al., 2021). The SDF represents the 3D distance from a query point to the nearest surface, implicitly capturing the 3D geometry.
Given a feature volume $\mathcal{V}$, we apply a shallow 3D CNN $\phi$ to directly produce three outputs: an SDF feature volume $\mathcal{S}\in\mathbb{R}^{X\times Y\times Z}$, a spherical harmonic (SH) (Yu et al., 2021; Zhu et al., 2023a) coefficient field $\mathcal{K}\in\mathbb{R}^{D\times X\times Y\times Z}$ (where $D=3\cdot(l_{\mathrm{max}}+1)^{2}$) for color rendering, and a semantic feature volume $\mathcal{F}\in\mathbb{R}^{C_{\mathrm{semantic}}\times X\times Y\times Z}$:
$\mathcal{S}\in\mathbb{R}^{X\times Y\times Z},\quad\mathcal{K}\in\mathbb{R}^{D\times X\times Y\times Z},\quad\mathcal{F}\in\mathbb{R}^{C_{\mathrm{semantic}}\times X\times Y\times Z}=\phi(\mathcal{V}).$ (2)
Unlike prior work (Huang et al., 2023; Zhu et al., 2023b; Yang et al., 2024a), which employs an MLP to compute the attributes of each sampled point individually, we directly apply a 3D CNN to $\mathcal{V}$. This eliminates the need for pointwise MLP computations, reducing redundant processing and enabling more efficient execution. Consequently, our approach leads to substantial improvements in both time and memory efficiency, especially when sampling a large number of points during rendering.
To render a 2D pixel $i$, we sample $N$ ray points $\{\mathbf{p}_{j}=\mathbf{o}+t_{j}\mathbf{d}_{i}\mid j=1,\ldots,N,\,t_{j}<t_{j+1}\}$ from ray $\mathbf{r}_{i}$, where $\mathbf{o}$ is the camera origin and $\mathbf{d}_{i}$ is the viewing direction. Attributes for each point are obtained via trilinear sampling:
$s_{j}=\tau(\mathcal{S},\mathbf{p}_{j}),\quad\mathbf{k}_{j}=\tau(\mathcal{K},\mathbf{p}_{j}),\quad\mathbf{f}_{j}=\tau(\mathcal{F},\mathbf{p}_{j}).$ (3)
The SH vector $\mathbf{k}_{j}=(k_{l}^{m})_{0\leq l\leq l_{\text{max}},-l\leq m\leq l}$, where $k_{l}^{m}\in\mathbb{R}^{3}$, is used to compute view-dependent colors $\hat{\mathbf{c}}_{j}$ by querying the SH basis functions $Y_{l}^{m}:\mathbb{S}^{2}\rightarrow\mathbb{R}$ based on the viewing direction $\mathbf{d}_{j}$:
$\hat{\mathbf{c}}_{j}=\operatorname{Sigmoid}\mathopen{}\mathclose{{}\left(\sum_{l=0}^{l_{\text{max}}}\sum_{m=-l}^{l}k_{l}^{m}Y_{l}^{m}(\mathbf{d}_{j})}\right).$ (4)
Following the formulation in NeuS (Wang et al., 2021), the RGB color $\hat{\mathbf{C}}_{i}$, depth $\hat{\mathbf{D}}_{i}$, and semantic feature $\hat{\mathbf{F}}_{i}$ for pixel $i$ are computed by integrating the predicted values along the ray:
$\hat{\mathbf{C}}_{i}=\sum_{j=1}^{N}w_{j}\hat{\mathbf{c}}_{j},\quad\hat{\mathbf{D}}_{i}=\sum_{j=1}^{N}w_{j}t_{j},\quad\hat{\mathbf{F}}_{i}=\sum_{j=1}^{N}w_{j}\hat{\mathbf{f}}_{j},$ (5)
where $w_{j}=T_{j}\alpha_{j}$ is the occlusion-aware weight, with $T_{j}=\prod_{k=1}^{j-1}(1-\alpha_{k})$ representing the accumulated transmittance and $\alpha_{j}$ being the opacity value. Specifically, $\alpha_{j}$ is computed as:
$\alpha_{j}=\max\mathopen{}\mathclose{{}\left(\frac{\sigma_{s}(s_{j})-\sigma_{s}(s_{j+1})}{\sigma_{s}(s_{j})},0}\right),$ (6)
where $\sigma_{s}(x)=(1+e^{-sx})^{-1}$ is the sigmoid function modulated by a learnable parameter $s$.
2.4 Loss Functions
During pre-training, we randomly sample $K$ pixels from multi-view inputs in each iteration. The rendering loss is calculated based on the differences between the input pixel values and the predicted values. For the semantic feature map, we use the feature map from AM-RADIO (Ranzinger et al., 2024) as supervision. Our framework has the capability to distill knowledge from multiple vision foundation models by adding multiple rendering heads. However, this paper does not explore that approach, as it is not the primary focus. The rendering loss is expressed as:
$\mathcal{L}_{\text{render}}=\frac{1}{K}\sum_{i=1}^{K}\mathopen{}\mathclose{{}\left(\lambda_{\text{color}}\cdot\|\mathbf{C}_{i}-\hat{\mathbf{C}}_{i}\|+\lambda_{\text{depth}}\cdot\|\mathbf{D}_{i}-\hat{\mathbf{D}}_{i}\|+\lambda_{\text{semantic}}\cdot\|\mathbf{F}_{i}-\hat{\mathbf{F}}_{i}\|}\right).$ (7)
Additionally, we incorporate the Eikonal regularization loss $\mathcal{L}_{\text{eikonal}}$, near-surface SDF supervision loss $\mathcal{L}_{\text{sdf}}$, and free space SDF loss $\mathcal{L}_{\text{free}}$, which are standard in neural surface reconstruction. Detailed definitions of these losses are provided in Appendix A. The total loss is defined as:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{render}}+\lambda_{\text{eikonal}}\cdot\mathcal{L}_{\text{eikonal}}+\lambda_{\text{sdf}}\cdot\mathcal{L}_{\text{sdf}}+\lambda_{\text{free}}\cdot\mathcal{L}_{\text{free}}.$ (8)
Figure 3: Overview of our large-scale embodied evaluation. We conduct the largest-scale evaluation of embodied representation learning to date. Our study encompasses 268 tasks across 8 simulators, including both single-task and language-conditioned multi-task settings. We evaluate diverse policy architectures and assess various state-of-the-art representation methods. This thorough evaluation allows us to provide a comprehensive and unbiased analysis of different representations.
3 Large-Scale Embodied Evaluation
Unlike the CV or NLP communities, where large-scale benchmarks are common, embodied representations have not been thoroughly assessed. The largest previous evaluation, VC-1 (Majumdar et al., 2023), includes only 17 tasks. This may lead to randomness and bias. Therefore, we have created the largest embodied evaluation to date, encompassing 268 tasks across 8 simulators—over 15 times larger than VC-1’s evaluation. Additionally, unlike previous approaches (Majumdar et al., 2023; Nair et al., 2022; Radosavovic et al., 2023) that used a small MLP policy under single-task settings, our evaluation spans multiple policy types (e.g. MLP, diffusion, transformer) and includes both single-task and language-conditioned multi-task settings. This unprecedented scale and diversity ensure robust and convincing conclusions. During all evaluations, we adhere to standard practices by freezing the pre-trained representation model. Our detailed evaluation settings can be found in Appendix B. The overview of our evaluation is shown in Fig. 3.
We have included 3 single-task benchmarks:
1) VC-1 (Majumdar et al., 2023) involves 4 selected simulators with 14 tasks in total: Adroit (AD) (Kumar, 2016), Meta-World (MW) (Yu et al., 2020), DMControl (DMC) (Tunyasuvunakool et al., 2020), and TriFinger (TF) (Wüthrich et al., 2020). We use a 3-layer MLP as the policy network.
2) Franka Kitchen (Gupta et al., 2019) involves 5 selected tasks. Each task spans two camera viewpoints and three random seeds. We utilize 25 demonstrations to train a 2-layer MLP policy.
3) Meta-World (Yu et al., 2020) involves 48 selected tasks of varying difficulty. We implemented the Diffusion Policy (Chi et al., 2023) on this benchmark and adhered to the setup in Ze et al. (2024) to generate 10 demonstrations for each single-task training, followed by evaluation through 20 rollouts.
We have also included 2 language-conditioned multi-task benchmarks:
1) RLBench (James et al., 2020) features 71 selected tasks that can be successfully executed. We divide the tasks into two groups according to their category defined by PolarNet (Chen et al., 2023). We employ RVT-2 (Goyal et al., 2024), the SOTA method on this benchmark, as our policy.
2) LIBERO (Liu et al., 2024) comprises 130 tasks across 5 suites: LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, LIBERO-10, and LIBERO-90. We train a language-conditioned transformer policy provided by the original LIBERO on each suite with only 20 demonstrations per task.
4 Training and Implementation Details
In this section, we present the implementation and training of our SPA model. We first compile several multi-view datasets, training ViT-B models on each to assess the impact of different datasets (Sec. 4.1). Finally, we integrate all factors and scale up both data and model size to train the strongest version of SPA using a ViT-large (ViT-L) backbone (Sec. 4.2). More details can be found in Appendix C.
4.1 Dataset Investigation
We collect several multi-view datasets. To investigate their effectiveness in SPA representation learning, we train a ViT-B model on one or two of the datasets, keeping the total training steps constant, and assess performance on the VC-1 benchmarks. For simplicity, semantic rendering is disabled. The datasets investigated are listed in the first column of Tab. 1. Most datasets provide ground-truth depth, which we use for supervision. As our findings above reveal that depth supervision is helpful, for datasets lacking ground-truth depth, we employ a depth estimation model. For instance, Droid (Khazatsky et al., 2024) only offers binocular images, so we apply CroCo-Stereo (Weinzaepfel et al., 2023) for dense depth estimation. Additionally, due to inaccurate camera poses in Droid, we treat its data as single-view inputs. The results are presented in Tab. 1, with further details in Appendix C. Our analysis reveals that some datasets can be detrimental. For example, although RH20T (Fang et al., 2023a) is a large-scale robotic dataset, its lack of visual diversity—stemming from data collected in the same lab—negatively impacts representation learning.
Table 1: Influence of different datasets. We present the performance results on the VC-1 benchmark. Mean S.R. refers to the mean success rate across all individual tasks.
Datasets AD MW DMC TF Mean S.R.
ScanNet (Dai et al., 2017) 52.67$\pm$4.11 90.93$\pm$3.22 65.11$\pm$1.31 70.75$\pm$1.08 73.68
ScanNet++ (Yeshwanth et al., 2023) 56.00$\pm$2.83 89.87$\pm$4.20 62.24$\pm$4.51 71.28$\pm$0.38 72.51
Arkitscenes (Baruch et al., 2021) 50.67$\pm$5.73 89.87$\pm$4.59 60.51$\pm$2.55 66.54$\pm$0.13 70.45
Droid (Khazatsky et al., 2024) 53.33$\pm$5.25 90.40$\pm$4.90 60.99$\pm$3.72 73.28$\pm$0.61 72.16
Hypersim (Roberts et al., 2021) 52.67$\pm$4.11 88.80$\pm$3.27 60.84$\pm$2.06 72.29$\pm$0.47 71.29
Hypersim + ADT (Pan et al., 2023) 52.00$\pm$2.83 87.20$\pm$2.30 63.61$\pm$1.04 70.83$\pm$0.13 71.41
Hypersim + S3DIS (Armeni et al., 2017) 49.33$\pm$0.94 94.13$\pm$2.04 64.57$\pm$3.91 71.74$\pm$0.75 73.98
Hypersim + Structured3D (Zheng et al., 2020) 46.67$\pm$4.11 80.27$\pm$7.72 58.02$\pm$ 2.34 65.05$\pm$0.40 65.35
Hypersim + RH20T (Fang et al., 2023a) 47.33$\pm$1.89 86.93$\pm$4.99 57.01$\pm$4.35 64.28$\pm$0.46 67.35
Hypersim + ASE (Avetisyan et al., 2024) 47.33$\pm$4.11 87.73$\pm$3.39 60.62$\pm$4.14 68.59$\pm$0.30 69.54
4.2 Put All Together
Based on the previous analyses, we proceed to pre-train the final version of SPA. We use a mask ratio of $0.5$ and enable all three rendering losses. Following Ponder (Huang et al., 2023), we set the weight for the RGB loss to $10$, the weights for the depth and semantic losses to $1$, and use $\lambda_{\mathrm{eikonal}}=0.01$, $\lambda_{\mathrm{sdf}}=10$, and $\lambda_{\mathrm{free}}=1$. The volume size is $128\times 128\times 32$. For stable training, we apply the Exponential Moving Average (EMA) technique with a decay of $0.999$. We use AdamW (Loshchilov et al., 2017) as the optimizer with a weight decay of $0.04$ and a learning rate of $8e^{-4}$. OneCycle (Smith & Topin, 2019) learning rate scheduler is adopted. We utilize 80 NVIDIA A100-SXM4-80GB GPUs, each with a batch size of $2$, and accumulate gradients over 8 batches, resulting in a total effective batch size of $2\times 8\times 80=1280$. Training is conducted over $2000$ epochs, sampling each dataset to match the size of ADT per epoch. The datasets used for the final version include ScanNet, ScanNet++, ADT, S3DIS, Hypersim, and Droid.
5 Experiment Results
In this section, we present the results of our large-scale evaluation. Our experiments are designed to address the following research questions:
Q1: How does SPA compare to other methods in our large-scale embodied evaluation?
Q2: What insights do we gain about various representation learning approaches from our evaluation?
Q3: Does SPA really learn enhanced 3D awareness that results in improved embodied representation?
Q4: Can SPA facilitate robot learning in real-world environments in a zero-shot manner?
5.1 Overall Comparisons (Q1, Q2)
Table 2: Summary of different representation learning methods. ‘#Param.’ is the total parameters of the encoder, while ‘#Frames’ indicates the total number of image frames used during pre-training.
Method Vision-Centric Multi-Modal Embodied-Specific Distilled AM-RADIO (Ranzinger et al., 2024)
MoCoV3 (Chen et al., 2020b) MAE (He et al., 2022) DINOV2 (Oquab et al., 2023) CLIP (Radford et al., 2021) EVA (Fang et al., 2023b) InternViT-300M (Chen et al., 2024b) InternViT-6B (Chen et al., 2024b) MVP (Radosavovic et al., 2023) VC-1 (Majumdar et al., 2023) SPA (Ours)
Is Vanilla?
Input Size 224 224 224 224 224 448 224 256 224 224 dynamic
Patch Size 16 16 14 14 14 14 14 16 16 16 16
#Param. 303M 303M 303M 303M 303M 303M 5.9B 303M 303M 303M 653M
#Frames 1.28M 1.28M 1.2B 400M 14M 5.0B 5.0B 4.5M 5.6M 3.8M 1.4B
Table 3: Comparison of different representation learning methods. ‘OOM’ indicates an out-of-memory error during evaluation. The best and second-best results are bolded and underlined respectively. The number in parentheses denotes the number of tasks. S.R. denotes ‘Success Rate’.
Method Vision-Centric Multi-Modal Embodied-Specific
Benchmark MoCoV3 MAE DINOV2 CLIP EVA InternViT- 300M InternViT- 6B MVP VC-1 SPA (Ours)
VC-1 AD (2) 58.7$\pm$7.0 58.0$\pm$2.0 47.3$\pm$3.1 48.7$\pm$3.1 58.0$\pm$6.0 53.3$\pm$3.1 60.0$\pm$9.2 53.3$\pm$4.2 54.0$\pm$4.0 60.0$\pm$4.0
MW (5) 88.8$\pm$5.0 90.0$\pm$4.6 84.0$\pm$3.7 77.1$\pm$3.2 90.7$\pm$0.9 84.0$\pm$3.7 89.1$\pm$1.2 93.6$\pm$5.2 87.5$\pm$3.8 93.3$\pm$2.0
DMC (5) 67.3$\pm$3.3 74.4$\pm$1.8 64.5$\pm$2.5 53.9$\pm$3.6 62.7$\pm$2.8 53.3$\pm$0.4 66.3$\pm$3.2 69.4$\pm$2.6 65.3$\pm$3.6 71.1$\pm$5.0
TF (2) 67.9$\pm$0.2 73.0$\pm$0.5 68.5$\pm$0.4 56.1$\pm$1.6 67.2$\pm$0.2 65.2$\pm$1.6 70.7$\pm$0.9 73.2$\pm$0.8 70.9$\pm$1.1 73.6$\pm$2.0
RLBench Group 1 (35) 73.7 78.3 78.2 76.8 75.2 74.1 OOM 76.2 80.1 80.5
Group 2 (36) 54.2 57.7 56.1 55.7 57.0 54.9 OOM 56.3 55.7 61.2
Meta-World (48) 69.3$\pm$1.5 67.8$\pm$1.7 56.3$\pm$0.6 66.7$\pm$1.7 63.7$\pm$1.3 57.5$\pm$1.7 OOM 66.4$\pm$1.7 68.6$\pm$1.5 69.2$\pm$1.7
LIBERO Object (10) 65.3$\pm$8.0 71.7$\pm$13.1 64.7$\pm$9.9 50.2$\pm$7.0 73.2$\pm$6.0 67.7$\pm$6.0 58.0$\pm$10.6 63.7$\pm$4.8 69.7$\pm$7.2 76.7$\pm$5.3
Spatial (10) 40.5$\pm$0.9 57.2$\pm$2.9 36.3$\pm$11.8 32.2$\pm$0.6 59.3$\pm$7.7 48.3$\pm$6.4 42.0$\pm$10.3 58.0$\pm$6.2 50.5$\pm$7.5 50.0$\pm$3.8
Goal (10) 49.2$\pm$8.1 54.3$\pm$6.0 22.2$\pm$2.3 30.3$\pm$3.2 56.8$\pm$2.9 58.8$\pm$4.5 33.2$\pm$2.0 63.8$\pm$2.8 57.5$\pm$6.6 65.3$\pm$2.5
10 (10) 34.2$\pm$3.8 41.2$\pm$4.5 28.3$\pm$3.0 27.5$\pm$3.9 43.3$\pm$2.8 38.2$\pm$1.3 34.3$\pm$4.6 39.0$\pm$0.9 39.7$\pm$3.5 40.2$\pm$3.6
90 (90) 30.0$\pm$1.4 29.9$\pm$2.0 27.5$\pm$2.2 29.4$\pm$2.0 31.3$\pm$2.3 23.8$\pm$1.8 27.1$\pm$2.1 32.1$\pm$3.5 30.6$\pm$3.3 32.2$\pm$1.6
Franka-Kitchen (5) 48.3$\pm$4.7 42.7$\pm$2.6 40.9$\pm$6.4 30.8$\pm$3.3 37.3$\pm$1.3 28.5$\pm$1.7 OOM 34.3$\pm$6.1 37.5$\pm$3.5 40.6$\pm$1.9
Mean S.R. $\uparrow$ 81.67 85.13 75.18 77.10 83.84 75.41 30.65 84.85 84.69 88.63
Mean Rank $\downarrow$ 4.51 4.07 5.61 5.17 4.37 5.92 7.57 4.24 4.13 3.20
Evaluation Metrics. We follow prior work (Majumdar et al., 2023; Zhu et al., 2024) in reporting two metrics: Mean Success Rate (Mean S.R.) and Mean Rank. Mean S.R. is the average success rate across all tasks, indicating overall performance, while Mean Rank reflects the average ranking of each method’s success rate across tasks, providing a measure of relative performance. Since RLBench has fixed train and test sets, we report a single result for this benchmark.
Baselines. We evaluate 9 SOTA representation learning models, all using ViT-L backbone, categorized into vision-centric, multi-modal, and embodied-specific. This also includes a 6B multi-modal model (Chen et al., 2024b). The vision-centric methods are originally from the vision community; the multi-modal methods are typically CLIP-style language-image pre-trained models and are used specifically for VLMs; the embodied-specific methods are designed and pre-trained specifically for embodied AI tasks. Details are summarized in Tab. 2. The results on each benchmark are shown in Tab. 3. For detailed results on each task and each random seed, please refer to Appendix D. We also have visualized the performance radar chart and the per-task rank distributions in Fig. 1.
Finding 1: We observe that SPA demonstrates superior performance in both mean success rate and mean rank. While no method ranks first across all individual benchmarks, consistent with the findings by Majumdar et al. (2023), SPA achieves the best or second-best mean success rate in 11 out of 13 benchmarks. Additionally, it ranks in the top 3 for over 65.5% of individual tasks, surpassing the second and third highest percentages of 46.8% for MAE and 46.0% for VC-1, respectively. These trends demonstrate the robustness and superiority of SPA.
Finding 2: We observe that for vision-centric methods, superior performance on vision tasks does not necessarily translate to better embodied performance. Despite using 10 times more data, DINOV2 performs worse than MoCoV3 and MAE. Notably, MAE performs exceptionally well, likely due to its reconstruction objective, which enhances 2D spatial awareness. Interestingly, methods like MVP and VC-1, which are MAE models pre-trained on human interaction data, show no clear advantage over ImageNet (Deng et al., 2009) pre-trained MAE. This suggests that while human activity data may seem more relevant, data diversity and thorough convergence are more critical.
Finding 3: Multimodal methods generally perform poorly in embodied evaluations, except EVA, which combines image-language contrastive techniques with MAE reconstruction. Furthermore, InternViT-6B, despite having significantly more model parameters, does not demonstrate superiority and even performs worse on some benchmarks compared to InternViT-300M. This indicates that current scaling properties of multimodal approaches do not effectively translate to embodied AI.
Finding 4: Focusing on a single benchmark can lead to highly biased conclusions. For instance, ImageNet pre-trained methods (e.g. MoCoV3 and MAE) perform exceptionally well on the Franka Kitchen benchmark, suggesting a minimal domain gap between ImageNet and Franka Kitchen observations. Moreover, despite being based on MAE, previous SOTA embodied representations like MVP and VC-1 do not consistently outperform the original ImageNet version. These observations underscore the importance of our large-scale embodied evaluation.
Table 4: Additional comparisons of ViT-base models. S.R. denotes ‘Success Rate’.
Methods DINOV2-B (Oquab et al., 2023) MAE-B (He et al., 2022) R3M-B (Nair et al., 2022) VC-1-B (Majumdar et al., 2023) STP-B (Yang et al., 2024b) Voltron-B (Karamcheti et al., 2023) Theia-B (Shang et al., 2024) SPA-B (Ours)
Is Vanilla?
Embodied?
VC-1 AD 36.67$\pm$2.31 52.67$\pm$3.06 48.00$\pm$6.93 50.00$\pm$5.29 52.00$\pm$2.00 46.67$\pm$4.62 53.33$\pm$5.03 52.00$\pm$3.46
MW 60.80$\pm$0.80 88.80$\pm$4.00 59.20$\pm$5.60 86.67$\pm$0.92 92.00$\pm$1.39 84.00$\pm$3.20 89.07$\pm$3.23 92.00$\pm$4.16
DMC 35.19$\pm$4.87 62.39$\pm$4.97 49.57$\pm$4.85 60.92$\pm$0.70 61.40$\pm$2.86 56.36$\pm$2.01 64.98$\pm$3.42 64.21$\pm$3.52
TF 54.50$\pm$1.16 70.78$\pm$0.17 56.18$\pm$7.00 72.33$\pm$0.69 67.96$\pm$0.95 74.26$\pm$1.57 69.41$\pm$0.60 73.06$\pm$0.51
Mean S.R. 47.31 71.63 54.37 70.19 71.92 69.50 72.55 73.66
5.2 Additional Comparisons (Q1)
We primarily compare with SOTA methods using the ViT-L backbone, which is commonly available and pre-trained on large-scale datasets. However, some embodied-specific models are only offered in ViT-B variants. Therefore, we provide additional comparisons with several ViT-B models in Tab. 4. Our ViT-B version, SPA-B, also outperforms other baselines. Furthermore, when compared to SPA-L on VC-1 benchmarks, the mean success rate increases by 4.16 (73.66 $\rightarrow$ 77.82). This indicates that increasing the model size positively impacts SPA’s performance.
5.3 Study on 3D Awareness of SPA (Q3)
Table 5: Zero-shot camera pose estimation. Trans. and Rot. denote ‘translation’ and ‘rotation’ errors respectively. The detailed metrics on the error calculation are listed in Appendix E.
Error MoCoV3 MAE DINOV2 CLIP EVA InternViT-300M InternViT-6B MVP VC-1 SPA(Ours)
Trans. (×$\times e^{-2}$) 2.29$\pm$0.07 2.15$\pm$0.07 6.55$\pm$0.07 4.21$\pm$0.37 5.49$\pm$0.24 4.62$\pm$0.14 5.39$\pm$0.41 2.15$\pm$0.12 2.02$\pm$0.07 1.65$\pm$0.09
Rot. (×$\times e^{-1}$) 0.79$\pm$0.07 0.73$\pm$0.03 2.12$\pm$0.25 1.52$\pm$0.08 1.83$\pm$0.09 1.83$\pm$0.08 1.91$\pm$0.12 0.77$\pm$0.05 0.72$\pm$0.01 0.61$\pm$0.01
Table 6: Additional ablations on VC-1.
Methods SPA-B SPA-MAE RADIO E-RADIO
VC-1 AD 52.00$\pm$3.46 55.33$\pm$3.06 55.33$\pm$3.06 56.67$\pm$2.31
MW 92.00$\pm$4.16 90.67$\pm$6.00 72.00$\pm$9.23 83.47$\pm$4.11
DMC 64.21$\pm$3.52 63.85$\pm$3.60 67.38$\pm$7.35 62.92$\pm$4.24
TF 73.06$\pm$0.51 70.14$\pm$0.98 71.75$\pm$0.14 68.44$\pm$1.19
Mean S. R. 73.66 73.11 67.93 70.16
Firstly, we aim to provide clear evidence that the performance improvements of SPA are due to its 3D awareness. To demonstrate this, we conducted two additional ablation studies on the VC-1 benchmarks: 1) To determine whether the performance gain is due to SPA’s pre-training objectives or the datasets used, we continue pre-training the ImageNet pre-trained MAE-B (the most competitive method besides SPA) on the same datasets used by SPA-B, referring to this model as SPA-MAE. Hyperparameters, including mask ratio and batch size, are kept at their default settings, and both the ImageNet pre-trained encoder and decoder weights are initially loaded. 2) Since SPA uses the feature map of RADIO for semantic rendering supervision, we also evaluate the original RADIO (653M parameters) and its efficient version, E-RADIO (391M parameters). Results are presented in Tab. 6.
Figure 4: Correlation between mean success rate and camera pose regression error.
Finding 5: The 3D-aware pre-training objective significantly enhances SPA’s performance. It surpasses the single-image naive MAE with the same data. Notably, SPA learns superior representations compared to its semantic rendering teacher by a substantial margin.
Moreover, we provide both quantitative and qualitative evidence to demonstrate that SPA has acquired 3D awareness. For qualitative analysis, we visualize the zero-shot feature maps on multiview images of different encoder outputs, as shown in Fig. 5. The images are taken from the unseen Arkitscenes dataset. For quantitative analysis, we evaluate the zero-shot 3D awareness of various methods using a camera pose estimation task on the NAVI dataset (Jampani et al., 2023).
Figure 5: Feature map visualization.
Specifically, given a pair of images from different viewpoints, we use a frozen encoder to extract features and concatenate them. A small MLP then regresses the relative camera pose and we report rotation and translation errors in Tab. 5. Details are in Appendix E. While El Banani et al. (2024) has explored 3D awareness of different vision models, their context differs. Their tasks can allow strong semantic models like DINOV2 to ‘cheat’. For example, multiview correspondence can be achieved through semantic matching, and the relative depth estimation task involves transforming normalized values into discrete bins, resembling a per-pixel classification task. Additionally, they emphasize fine-grained dense local context, whereas, embodied AI focuses more on sparse, global information (Nair et al., 2022). Thus, we believe camera pose estimation, which predicts a global ‘pose’ from observations, is more relevant to embodied AI, where a policy must predict a global ‘action’.
Finding 6: We observe that SPA outperforms all other methods in zero-shot camera pose estimation. It achieved an 18.3% improvement in translation and a 15.3% reduction in rotation error compared to the second-best model. Additionally, we identify a clear positive correlation between camera pose estimation and embodied evaluation performance, as demonstrated in Fig. 4. This finding supports our spatial hypothesis and may offer valuable insights for future research on embodied representation.
Finding 7: The feature map visualization provides clear evidence that SPA has learned multi-view consistent knowledge, demonstrating its 3D awareness. Additionally, the features produced by SPA are cleaner and more coherent. Though VC-1 also generates smooth features, they are not consistent across viewpoints. The feature maps from the multi-modal approach are highly noisy and lack details.
5.4 Hyperparameter Investigation
Table 7: Mask ratio and loss components. C., D., S. denote color, depth, and semantic.
Mask Ratio Loss VC-1 Benchmark Mean S.R.
C. D. S. AD MW DMC TF
0.00 53.3$\pm$4.6 88.5$\pm$5.7 57.5$\pm$2.6 74.1$\pm$0.6 70.36
0.25 52.7$\pm$3.1 89.6$\pm$4.5 57.6$\pm$3.0 70.4$\pm$1.7 70.17
0.50 53.3$\pm$4.2 88.8$\pm$1.6 60.1$\pm$3.1 72.6$\pm$0.7 71.18
0.75 51.3$\pm$1.2 88.0$\pm$3.5 61.1$\pm$3.5 73.0$\pm$0.8 71.01
0.95 51.3$\pm$1.2 85.6$\pm$4.0 62.5$\pm$5.3 73.1$\pm$0.2 70.67
0.50 51.3$\pm$1.2 90.9$\pm$3.3 58.8$\pm$5.6 71.5$\pm$1.0 71.01
0.50 52.0$\pm$2.0 89.3$\pm$3.3 53.9$\pm$4.3 70.9$\pm$1.3 68.71
0.50 52.7$\pm$3.1 88.0$\pm$4.5 61.5$\pm$3.4 71.6$\pm$1.2 71.16
We conduct hyperparameter tuning with a ViT-B model on ScanNet (Dai et al., 2017), and evaluate it on VC-1 benchmarks, as shown in Tab. 7. 1) Mask Ratio. Our results indicate that a mask ratio of $0.5$ is the most effective. 2) Loss Components. As discussed in Sec. 2.4, our rendering loss consists of color, depth, and semantic components. We sequentially deactivate each and find that all three are valuable. However, deactivating the semantic loss has the least impact.
5.5 Real-World Experiments (Q4)
Figure 6: Real-world task illustrations.
We conduct several real-world experiments to further investigate the generalization ability of different representations. Specifically, we utilize the open-sourced Low-Cost Robot Arm (Koch, 2024) to learn real-world tasks from pixels, with only 50 demonstrations per task using different frozen pre-trained representations. The robot performed two single-arm tasks: (1) picking a cube, and (2) stacking a yellow cube on a pink cube, as well as one dual-arm task: folding a cloth in half. Refer to Fig. 6 for illustrations and Appendix F for more details. We evaluate each task with 25 rollouts, with the results presented in Tab. 8. SPA consistently performs better on real-world tasks, suggesting that SPA’s pre-trained representations can robustly adapt to real-world environments without finetuning.
Table 8: Real-world experiment results. S.R. denotes ‘Success Rate’.
Methods MoCoV3 MAE DINOV2 CLIP EVA InternViT-300M InternViT-6B MVP VC-1 SPA (Ours)
Picking Cube 28.00 64.00 20.00 28.00 56.00 32.00 52.00 36.00 40.00 64.00
Stacking Cube 16.00 32.00 4.00 16.00 8.00 8.00 36.00 20.00 16.00 48.00
Folding Cloth 48.00 64.00 32.00 24.00 28.00 48.00 44.00 64.00 60.00 84.00
Mean S.R. 30.67 53.33 18.67 22.67 30.67 29.33 44.00 40.00 38.67 65.33
6 Related Work
Representation Learning for Computer Vision. Recent advances in computer vision have increasingly focused on unsupervised and self-supervised learning to utilize large amounts of unlabeled data. Techniques like contrastive learning (Chen et al., 2020a; 2021; b; He et al., 2020), masked autoencoders (He et al., 2022; Feichtenhofer et al., 2022; Bachmann et al., 2022; Tong et al., 2022; Wang et al., 2023), and self-distillation (Caron et al., 2021; Oquab et al., 2023; Ranzinger et al., 2024) have shown that effective representations can be learned without supervision. Moreover, multi-modal pre-training approaches (Radford et al., 2021; Fang et al., 2023b; Chen et al., 2024b) leverage language to learn more comprehensive representations. These developments have significantly improved transfer learning capabilities while also displaying zero-shot abilities.
Representation Learning for Embodied AI. Recent advances in embodied AI representation learning, inspired by computer vision, have applied techniques such as contrastive (Nair et al., 2022; Yang et al., 2023) and masked autoencoders (Radosavovic et al., 2023; Majumdar et al., 2023; Karamcheti et al., 2023; Yang et al., 2024b) to embodied AI. However, these approaches often emphasize semantic learning while overlooking the specific needs of embodied AI tasks. In this work, we propose a spatial hypothesis specifically for embodied AI representation learning, and we demonstrate how a standard 2D backbone can integrate 3D spatial awareness.
3D Robot Learning and 3D-Aware Computer Vision. Prior work in 3D robot learning has often relied on explicit 3D input (Zhu et al., 2024; Ze et al., 2024; Wang et al., 2024b; a; Shridhar et al., 2023; Chen et al., 2023), or lifting 2D features into 3D spaces (Ke et al., 2024; Goyal et al., 2024), providing a strong foundation for our spatial hypothesis. Given the scalability challenges of explicit 3D observations, some computer vision research has explored integrating 3D spatial awareness into 2D backbones (Yang et al., 2024a; Zhu et al., 2023b; Yue et al., 2025; Zhang et al., 2024). To the best of our knowledge, SPA is the first to systematically investigate this approach in embodied AI.
Neural Rendering. Recent advances in 3D vision, particularly in neural rendering (Mildenhall et al., 2021), have enabled the encoding of scenes using neural networks, which support differentiable rendering and reconstruction. Alongside improvements in neural rendering techniques themselves (Wang et al., 2021; Zhu et al., 2023a; Gropp et al., 2020; Ortiz et al., 2022; Wang et al., 2022), the Ponder series (Huang et al., 2023; Zhu et al., 2023b; Yang et al., 2024a) and subsequent works (Wang et al., 2024c; Irshad et al., 2024) have applied differentiable neural rendering for representation learning. However, they have primarily focused on 3D perception and autonomous driving scenarios. To the best of our knowledge, our work is the first to apply neural rendering for embodied AI representation learning using a standard 2D backbone, marking a novel contribution to this area of research.
7 Conclusion, Limitations, and Future Work
In this work, we propose that 3D spatial awareness is crucial for embodied AI and introduce SPA, a novel framework that pre-trains a standard ViT backbone with 3D spatial awareness. To validate our hypothesis, we conduct the largest-scale embodied evaluation to date, over 15 times larger than previous studies. Our experiments demonstrate the clear superiority of SPA and highlight the importance of 3D awareness. Despite strong results across simulated and real robotic tasks, limitations remain. Our evaluation is currently restricted to imitation learning (specifically behavior cloning), and exploring SPA’s performance in other settings, such as reinforcement learning, presents an exciting future direction. Incorporating SPA into VLMs to enhance their performance on spatial aware tasks (Chen et al., 2024a; Majumdar et al., 2024) also represents an exciting research area. Additionally, SPA currently focuses on static multi-view scenes; extending it to dynamic, temporal scenarios could enhance its generality. Lastly, while we use the ViT encoder for fair comparison, the volume decoder’s multi-view interaction knowledge could be leveraged in policy learning, offering further potential for improvement.
Acknowledgments
This work is supported by the National Key R&D Program of China (NO.2022ZD0160102) and Shanghai Artificial Intelligence Laboratory.
Appendix A Additional Rendering Losses
Here we detail the three additional rendering losses we have applied in Sec. 2.4.
Eikonal Regularization Loss. The Eikonal regularization loss, denoted as $\mathcal{L}_{\text{eikonal}}$, is a widely used loss function for the regularization of signed distance functions (SDFs) (Gropp et al., 2020). It is defined as:
$\mathcal{L}_{\text{eikonal}}=\frac{1}{N_{r}N_{p}}\sum_{i=1}^{N_{r}}\sum_{j=1}^{N_{p}}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left\|\nabla s(\mathbf{p}_{i,j})}\right\|-1}\right)^{2},$ (9)
where $\nabla s(\mathbf{p}_{i,j})$ represents the gradient of the SDF $s$ at the location $\mathbf{p}_{i,j}$. Since the SDF is a distance measure, $\mathcal{L}_{\text{eikonal}}$ encourages the gradients to have unit norm at the query point.
Near-Surface and Free Space Loss for SDF. To improve SDF estimation, we incorporate additional approximate SDF supervision, similar to iSDF (Ortiz et al., 2022) and GO-Surf (Wang et al., 2022). Specifically, for near-surface points, the difference between rendered depth and ground-truth depth serves as pseudo-SDF ground-truth supervision. For points far from the surface, a free space loss is used to further regularize the SDF values.
To compute the approximate SDF supervision, we define an indicator $b(z)$ for each sampled ray point with ray length $z$ and corresponding ground-truth depth $D$:
$b(z)=D-z.$ (10)
The value $b(z)$ can be considered a credible approximate SDF value when it is small. Let $t$ be a user-defined threshold, set to $0.05$ in our experiments. For sampled ray points satisfying $b(z)\leq t$, we apply the near-surface SDF loss to constrain the SDF prediction $s(z_{i,j})$:
$\mathcal{L}_{\text{sdf}}=\frac{1}{N_{r}N_{p}}\sum_{i=1}^{N_{r}}\sum_{j=1}^{N_{p}}\mathopen{}\mathclose{{}\left|s(z_{i,j})-b(z_{i,j})}\right|.$ (11)
For the remaining sampled ray points, we utilize a free space loss:
$\mathcal{L}_{\text{free}}=\frac{1}{N_{r}N_{p}}\sum_{i=1}^{N_{r}}\sum_{j=1}^{N_{p}}\max\mathopen{}\mathclose{{}\left(0,e^{-\alpha\cdot s(z_{i,j})}-1,s(z_{i,j})-b(z_{i,j})}\right),$ (12)
where $\alpha$ is set to 5, following Ortiz et al. (2022); Wang et al. (2022). Due to the presence of noisy depth images, $\mathcal{L}_{\text{sdf}}$ and $\mathcal{L}_{\text{free}}$ are applied only to rays with valid depth values.
In our experiments, we adopt a similar weighting scheme to GO-Surf (Wang et al., 2022), setting $\lambda_{C}=10.0$, $\lambda_{D}=1.0$, $\lambda_{\text{sdf}}=10.0$, and $\lambda_{\text{free}}=1.0$. We observe that the Eikonal term can lead to overly smooth reconstructions, so we use a small weight of $0.01$ for the Eikonal loss.
Appendix B Evaluation Setups
Here we detail the setups of our large-scale evaluation in Sec. 3. For the detailed visualizations of each task, we recommend the readers to read the original simulator’s or benchmark’s dataset.
B.1 Single-Task Benchmarks
VC-1 (Majumdar et al., 2023). This benchmark includes several simulators. We selected four: Adroit (Kumar, 2016), Meta-World (Yu et al., 2020), DMControl (Tunyasuvunakool et al., 2020), and TriFinger (Wüthrich et al., 2020). The Adroit subset focuses on dexterous manipulation with 2 tasks: Relocate and Pen. The Meta-World subset addresses two-finger gripper manipulation with 5 tasks: Button Press Topdown, Drawer Open, Bin Picking, Hammer, and Assembly. The DMControl subset is for locomotion control, also with 5 tasks: Walker Stand, Walker Walk, Reacher Easy, Cheetah Run, and Finger Spin. The TriFinger subset targets three-finger manipulation with 2 tasks: Reach Cube and Move Cube. For all tasks, we use a 3-layer MLP as the policy network for each single-task training, following the original implementation. Each task is trained with 100 demonstrations, except for 25 on Meta-World, and evaluated 50 times using the specific seeds 100, 200, and 300. The [CLS] token of a frozen pre-trained ViT is used as the observation feature. All hyper-parameters are kept the same with the original implementation.
Franka Kitchen (Gupta et al., 2019). Franka-Kitchen is a MuJoCo-modeled simulation environment with a Franka robot in a kitchen scene. Its action space is the 9-dimensional joint velocity with 7 DoF for the arm and 2 DoF for the gripper. Following previous works (Nair et al., 2022; Karamcheti et al., 2023), we evaluate five tasks: Sliding Door, Turning Light On, Opening Door, Turning Knob, and Opening Microwave. Each task spans two camera viewpoints and three random seeds. Similar to the evaluation scheme in VC-1, we utilize 25 demonstrations to train a policy model, which is a 2-layer MLP with hidden sizes [256, 256] preceded by a BatchNorm.
Meta-World (Yu et al., 2020). This benchmark comprises a series of tasks in which an agent directs a Sawyer robot arm to manipulate objects in a tabletop environment. We selected 48 tasks, encompassing easy, medium, and hard levels. We implemented the Diffusion Policy (Chi et al., 2023) on this benchmark and adhered to the setup in Ze et al. (2024) to generate 10 demonstrations for each single-task training, followed by evaluation through 20 rollouts. The average results across three fixed seeds (100, 200, 300) are reported. The [CLS] token from a frozen pre-trained ViT serves as the observation feature. The 48 tasks include: Button Press Wall, Door Close, Door Unlock, Drawer Close, Drawer Open, Faucet Close, Plate Slide, Plate Slide Back, Plate Slide Side, Window Close, Basketball, Bin Picking, Box Close, Coffee Push, Assembly, Disassemble, Push Wall, Shelf Place, Door Open, Button Press, Sweep Into, Door Lock, Reach Wall, Hammer, Stick Push, Button Press Topdown, Handle Press Side, Plate Slide Back Side, Sweep, Button Press Topdown Wall, Handle Press, Push, Coffee Pull, Dial Turn, Reach, Coffee Button, Pick Place Wall, Stick Pull, Hand Insert, Peg Insert Side, Pick Place, Faucet Open, Push Back, Lever Pull, Handle Pull, Soccer, Window Open, and Pick Out Of Hole.
B.2 Language-Conditioned Multi-Task Benchmarks
RLBench (James et al., 2020). This benchmark is a prominent language-conditioned multi-task robot learning framework. PolarNet (Chen et al., 2023) has categorized all tasks into 9 groups. We selected 71 tasks from RLBench that can be successfully executed and split them into two groups uniformly on categories: Group 1 with 35 tasks and Group 2 with 36 tasks. Each task includes 100 training demonstrations and 25 testing rollouts. For each group, we train a language-conditioned multi-task agent. We employ RVT-2 (Goyal et al., 2024), the state-of-the-art (SOTA) method on this benchmark, as our policy. RVT-2 takes multiple images rendered from point clouds as inputs and uses a convolutional block to generate feature maps. We substitute the convolutional block with different pre-trained ViTs, unpatchifying the latent vectors concatenated with the global [CLS] token to obtain feature maps. All other architectures and hyperparameters remain consistent with the original RVT-2 implementation.
The 35 tasks in Group 1 include: Basketball In Hoop, Put Rubbish In Bin, Meat Off Grill, Meat On Grill, Slide Block To Target, Reach And Drag, Take Frame Off Hanger, Water Plants, Hang Frame On Hanger, Wipe Desk, Stack Blocks, Reach Target, Push Button, Lamp On, Toilet Seat Down, Close Laptop Lid, Open Box, Open Drawer, Pick Up Cup, Turn Tap, Take Usb Out Of Computer, Play Jenga, Insert Onto Square Peg, Take Umbrella Out Of Umbrella Stand, Insert Usb In Computer, Straighten Rope, Turn Oven On, Change Clock, Close Microwave, Close Fridge, Close Grill, Open Grill, Unplug Charger, Press Switch, and Take Money Out Safe.
The 36 tasks in Group 2 include: Change Channel, Tv On, Push Buttons, Stack Wine, Scoop With Spatula, Place Hanger On Rack, Move Hanger, Sweep To Dustpan, Take Plate Off Colored Dish Rack, Screw Nail, Take Shoes Out Of Box, Slide Cabinet Open And Place Cups, Lamp Off, Pick And Lift, Take Lid Off Saucepan, Close Drawer, Close Box, Phone On Base, Toilet Seat Up, Put Books On Bookshelf, Beat The Buzz, Stack Cups, Put Knife On Chopping Board, Place Shape In Shape Sorter, Take Toilet Roll Off Stand, Put Umbrella In Umbrella Stand, Setup Checkers, Open Window, Open Wine Bottle, Open Microwave, Put Money In Safe, Open Door, Close Door, Open Fridge, Open Oven, and Plug Charger In Power Supply.
LIBERO (Liu et al., 2024). Built upon Robosuite (Zhu et al., 2020), LIBERO (Liu et al., 2024) generates a total of 130 language-conditioned tasks across five suites: LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, LIBERO-10, and LIBERO-90. Each suite contains 10 tasks, except for LIBERO-90, which includes 90 tasks. We train a language-conditioned multi-task policy for each suite, adopting the transformer policy provided by LIBERO. The image encoders are modified from default CNNs to frozen pre-trained ViTs, utilizing the [CLS] token for feature extraction. To expedite policy training, we use only 20 demonstrations per task and forgo augmentations, allowing for pre-extraction of all image features during training. After training for 25 epochs, the checkpoints from the 20th and 25th are evaluated with 20 rollouts per task, and the best checkpoint’s performance is taken. Finally, the results are averaged on 3 random seeds.
The 10 tasks in LIBERO-Spatial are:
1. Pick up the black bowl between the plate and the ramekin and place it on the plate.
2. Pick up the black bowl next to the ramekin and place it on the plate.
3. Pick up the black bowl from table center and place it on the plate.
4. Pick up the black bowl on the cookie box and place it on the plate.
5. Pick up the black bowl in the top drawer of the wooden cabinet and place it on the plate.
6. Pick up the black bowl on the ramekin and place it on the plate.
7. Pick up the black bowl next to the cookie box and place it on the plate.
8. Pick up the black bowl on the stove and place it on the plate.
9. Pick up the black bowl next to the plate and place it on the plate.
10. Pick up the black bowl on the wooden cabinet and place it on the plate.
The 10 tasks in LIBERO-Object are:
1. Pick up the alphabet soup and place it in the basket.
2. Pick up the cream cheese and place it in the basket.
3. Pick up the salad dressing and place it in the basket.
4. Pick up the BBQ sauce and place it in the basket.
5. Pick up the ketchup and place it in the basket.
6. Pick up the tomato sauce and place it in the basket.
7. Pick up the butter and place it in the basket.
8. Pick up the milk and place it in the basket.
9. Pick up the chocolate pudding and place it in the basket.
10. Pick up the orange juice and place it in the basket.
The 10 tasks in LIBERO-Goal are:
1. Open the middle drawer of the cabinet.
2. Put the bowl on the stove.
3. Put the wine bottle on top of the cabinet.
4. Open the top drawer and put the bowl inside.
5. Put the bowl on top of the cabinet.
6. Push the plate to the front of the stove.
7. Put the cream cheese in the bowl.
8. Turn on the stove.
9. Put the bowl on the plate.
10. Put the wine bottle on the rack.
The 10 tasks in LIBERO-10 are:
1. LIVING ROOM SCENE2: Put both the alphabet soup and the tomato sauce in the basket.
2. LIVING ROOM SCENE2: Put both the cream cheese box and the butter in the basket.
3. KITCHEN SCENE3: Turn on the stove and put the moka pot on it.
4. KITCHEN SCENE4: Put the black bowl in the bottom drawer of the cabinet and close it.
5. LIVING ROOM SCENE5: Put the white mug on the left plate and put the yellow and white mug on the right plate.
6. STUDY SCENE1: Pick up the book and place it in the back compartment of the caddy.
7. LIVING ROOM SCENE6: Put the white mug on the plate and put the chocolate pudding to the right of the plate.
8. LIVING ROOM SCENE1: Put both the alphabet soup and the cream cheese box in the basket.
9. KITCHEN SCENE8: Put both moka pots on the stove.
10. KITCHEN SCENE6: Put the yellow and white mug in the microwave and close it.
The 10 tasks in LIBERO-90 are:
1. KITCHEN SCENE10: Close the top drawer of the cabinet.
2. KITCHEN SCENE10: Close the top drawer of the cabinet and put the black bowl on top of it.
3. KITCHEN SCENE10: Put the black bowl in the top drawer of the cabinet.
4. KITCHEN SCENE10: Put the butter at the back in the top drawer of the cabinet and close it.
5. KITCHEN SCENE10: Put the butter at the front in the top drawer of the cabinet and close it.
6. KITCHEN SCENE10: Put the chocolate pudding in the top drawer of the cabinet and close it.
7. KITCHEN SCENE1: Open the bottom drawer of the cabinet.
8. KITCHEN SCENE1: Open the top drawer of the cabinet.
9. KITCHEN SCENE1: Open the top drawer of the cabinet and put the bowl in it.
10. KITCHEN SCENE1: Put the black bowl on the plate.
11. KITCHEN SCENE1: Put the black bowl on top of the cabinet.
12. KITCHEN SCENE2: Open the top drawer of the cabinet.
13. KITCHEN SCENE2: Put the black bowl at the back on the plate.
14. KITCHEN SCENE2: Put the black bowl at the front on the plate.
15. KITCHEN SCENE2: Put the middle black bowl on the plate.
16. KITCHEN SCENE2: Put the middle black bowl on top of the cabinet.
17. KITCHEN SCENE2: Stack the black bowl at the front on the black bowl in the middle.
18. KITCHEN SCENE2: Stack the middle black bowl on the back black bowl.
19. KITCHEN SCENE3: Put the frying pan on the stove.
20. KITCHEN SCENE3: Put the moka pot on the stove.
21. KITCHEN SCENE3: Turn on the stove.
22. KITCHEN SCENE3: Turn on the stove and put the frying pan on it.
23. KITCHEN SCENE4: Close the bottom drawer of the cabinet.
24. KITCHEN SCENE4: Close the bottom drawer of the cabinet and open the top drawer.
25. KITCHEN SCENE4: Put the black bowl in the bottom drawer of the cabinet.
26. KITCHEN SCENE4: Put the black bowl on top of the cabinet.
27. KITCHEN SCENE4: Put the wine bottle in the bottom drawer of the cabinet.
28. KITCHEN SCENE4: Put the wine bottle on the wine rack.
29. KITCHEN SCENE5: Close the top drawer of the cabinet.
30. KITCHEN SCENE5: Put the black bowl in the top drawer of the cabinet.
31. KITCHEN SCENE5: Put the black bowl on the plate.
32. KITCHEN SCENE5: Put the black bowl on top of the cabinet.
33. KITCHEN SCENE5: Put the ketchup in the top drawer of the cabinet.
34. KITCHEN SCENE6: Close the microwave.
35. KITCHEN SCENE6: Put the yellow and white mug to the front of the white mug.
36. KITCHEN SCENE7: Open the microwave.
37. KITCHEN SCENE7: Put the white bowl on the plate.
38. KITCHEN SCENE7: Put the white bowl to the right of the plate.
39. KITCHEN SCENE8: Put the right moka pot on the stove.
40. KITCHEN SCENE8: Turn off the stove.
41. KITCHEN SCENE9: Put the frying pan on the cabinet shelf.
42. KITCHEN SCENE9: Put the frying pan on top of the cabinet.
43. KITCHEN SCENE9: Put the frying pan under the cabinet shelf.
44. KITCHEN SCENE9: Put the white bowl on top of the cabinet.
45. KITCHEN SCENE9: Turn on the stove.
46. KITCHEN SCENE9: Turn on the stove and put the frying pan on it.
47. LIVING ROOM SCENE1: Pick up the alphabet soup and put it in the basket.
48. LIVING ROOM SCENE1: Pick up the cream cheese box and put it in the basket.
49. LIVING ROOM SCENE1: Pick up the ketchup and put it in the basket.
50. LIVING ROOM SCENE1: Pick up the tomato sauce and put it in the basket.
51. LIVING ROOM SCENE2: Pick up the alphabet soup and put it in the basket.
52. LIVING ROOM SCENE2: Pick up the butter and put it in the basket.
53. LIVING ROOM SCENE2: Pick up the milk and put it in the basket.
54. LIVING ROOM SCENE2: Pick up the orange juice and put it in the basket.
55. LIVING ROOM SCENE2: Pick up the tomato sauce and put it in the basket.
56. LIVING ROOM SCENE3: Pick up the alphabet soup and put it in the tray.
57. LIVING ROOM SCENE3: Pick up the butter and put it in the tray.
58. LIVING ROOM SCENE3: Pick up the cream cheese and put it in the tray.
59. LIVING ROOM SCENE3: Pick up the ketchup and put it in the tray.
60. LIVING ROOM SCENE3: Pick up the tomato sauce and put it in the tray.
61. LIVING ROOM SCENE4: Pick up the black bowl on the left and put it in the tray.
62. LIVING ROOM SCENE4: Pick up the chocolate pudding and put it in the tray.
63. LIVING ROOM SCENE4: Pick up the salad dressing and put it in the tray.
64. LIVING ROOM SCENE4: Stack the left bowl on the right bowl and place them in the tray.
65. LIVING ROOM SCENE4: Stack the right bowl on the left bowl and place them in the tray.
66. LIVING ROOM SCENE5: Put the red mug on the left plate.
67. LIVING ROOM SCENE5: Put the red mug on the right plate.
68. LIVING ROOM SCENE5: Put the white mug on the left plate.
69. LIVING ROOM SCENE5: Put the yellow and white mug on the right plate.
70. LIVING ROOM SCENE6: Put the chocolate pudding to the left of the plate.
71. LIVING ROOM SCENE6: Put the chocolate pudding to the right of the plate.
72. LIVING ROOM SCENE6: Put the red mug on the plate.
73. LIVING ROOM SCENE6: Put the white mug on the plate.
74. STUDY SCENE1: Pick up the book and place it in the front compartment of the caddy.
75. STUDY SCENE1: Pick up the book and place it in the left compartment of the caddy.
76. STUDY SCENE1: Pick up the book and place it in the right compartment of the caddy.
77. STUDY SCENE1: Pick up the yellow and white mug and place it to the right of the caddy.
78. STUDY SCENE2: Pick up the book and place it in the back compartment of the caddy.
79. STUDY SCENE2: Pick up the book and place it in the front compartment of the caddy.
80. STUDY SCENE2: Pick up the book and place it in the left compartment of the caddy.
81. STUDY SCENE2: Pick up the book and place it in the right compartment of the caddy.
82. STUDY SCENE3: Pick up the book and place it in the front compartment of the caddy.
83. STUDY SCENE3: Pick up the book and place it in the left compartment of the caddy.
84. STUDY SCENE3: Pick up the book and place it in the right compartment of the caddy.
85. STUDY SCENE3: Pick up the red mug and place it to the right of the caddy.
86. STUDY SCENE3: Pick up the white mug and place it to the right of the caddy.
87. STUDY SCENE4: Pick up the book in the middle and place it on the cabinet shelf.
88. STUDY SCENE4: Pick up the book on the left and place it on top of the shelf.
89. STUDY SCENE4: Pick up the book on the right and place it on the cabinet shelf.
90. STUDY SCENE4: Pick up the book on the right and place it under the cabinet shelf.
Appendix C More Implementation Details
C.1 Dataset Details
The datasets used for SPA include ScanNet, ScanNet++, Hypersim, ADT, S3DIS, and Droid.
ScanNet consists of 1.89 million frames in total. Each epoch includes 1.5 times the dataset size. For each scene, a random starting frame is selected, followed by the sampling of 1 to 8 frames at random, with an interval of 8 frames between them.
ScanNet++ comprises 0.11 million frames. Each epoch includes 5 times the dataset size. For each scene, a random starting frame is selected, followed by the sampling of 1 to 8 frames at random, with an interval of 5 frames between them.
Hypersim contains 0.03 million frames. Each epoch includes 8 times the dataset size. For each scene, we randomly select 1 to 8 continuous frames.
ADT consists of 0.0015M frames in total. Each epoch includes 1 times the dataset size. For each scene, 1 to 8 continuous frames are randomly selected.
S3DIS consists of 0.015 million frames. Each epoch includes 5 times the dataset size. For each scene, a random starting frame is selected, followed by the sampling of 1 to 8 frames at random, with an interval of 5 frames between them.
Droid contains a large number of videos, but due to the high similarity between frames, the videos are first downsampled by a factor of 15 during pre-processing, resulting in 1.78 million frames. Since Droid does not provide depth data, we utilize Croco-Stereo Weinzaepfel et al. (2023) to estimate dense depth maps for rendering supervision. Additionally, due to the significant noise in the camera pose data, only a single frame is sampled at a time during training.
During pre-training, we first resize the multi-view input images to slightly larger than $224\times 224$, and then randomly crop them to a final size of $224\times 224$. Random photometric distortions with a probability of $0.5$ are applied for augmentation, including brightness ranging from 0.875 to 1.125, contrast ranging from 0.5 to 1.5, saturation ranging from 0.5 to 1.5, and hue ranging from -0.05 to 0.05. Frames with very small valid depth areas or scene boxes are filtered out.
For semantic rendering supervision, we observe that using larger image sizes improves the quality of feature maps generated by RADIO. Consequently, we resize the images to $1024\times 1024$ before feeding them into RADIO, which outputs a feature map of size $64\times 64$. We then apply bilinear sampling to query the semantic feature labels for each pixel.
C.2 Pre-Training Details
For stability during pre-training, we apply the Exponential Moving Average (EMA) with a decay rate of 0.999. The model is trained for 2000 epochs on 80 NVIDIA A100-80G GPUs, using a gradient clipping threshold of 1.0. Each GPU processes a batch size of 2, with 8 gradient accumulation steps, resulting in a total effective batch size of $2\times 80\times 8=1280$. We employ the AdamW optimizer with a weight decay of 0.04. The base learning rate is set to $5\times 10^{-6}$, and the actual learning rate is scaled by a factor of 8 times the effective batch size. A OneCycle learning rate scheduler is used, with a percentage start of 0.05, a divide factor of 100, and a final divide factor of 1000.
To facilitate faster convergence and improve stability, we initialize the encoder with ImageNet pre-trained weights from the Masked Autoencoder (MAE), applying a learning rate layer decay of 0.8. This initialization does not affect the validity of our conclusions, as demonstrated by the ablation study of SPA-MAE in Sec. 5.3. The ViT encoder and upsampling layers are trained with FP16 precision, while the volume decoder is trained with FP32 precision.
We set the loss weights to $\lambda_{\text{color}}=10$, $\lambda_{\text{depth}}=1$, $\lambda_{\text{semantic}}=1$, $\lambda_{\text{eikonal}}=0.01$, $\lambda_{\text{free}}=1$, and $\lambda_{\text{sdf}}=10$. For the NeuS sampler, the initial number of samples is set to 72, with 24 importance samples. In each iteration, we randomly sample $512$ pixels per view for rendering and supervision.
Appendix D Detailed Results of Each Task
We present the results of all individual tasks in Tab. 12, Tab. 13, Tab. 14, Tab. 15, Tab. 16, and Tab. 17.
Appendix E Camera Pose Estimation Details
We adopt a setup similar to that of El Banani et al. (2024) for camera pose estimation using the NAVI dataset (Jampani et al., 2023). Given an image pair from different viewpoints, we first extract features from each image using a frozen, pre-trained Vision Transformer (ViT) encoder. Following standard protocols for embodied evaluation, we use the [CLS] token as the feature representation. The two [CLS] tokens are then concatenated and passed through a BatchNorm layer and a Multi-Layer Perceptron (MLP) to regress the camera pose. The MLP consists of four linear layers with three ReLU activations, using hidden sizes of 512, 256, and 128 units, and outputs a 7-dimensional pose vector. The first three dimensions represent the $xyz$ translation, while the last four dimensions correspond to the rotation quaternions.
We employ the Mean Squared Error (MSE) loss function and optimize the model using the AdamW optimizer with a OneCycle learning rate scheduler. The model is trained for 100 epochs with a base learning rate of $1\times 10^{-3}$ and a starting percentage of 0.1. For evaluation, we use Euclidean distance as the translation error metric and geodesic distance as the rotation error metric. The geodesic distance between two quaternions $q_{1}$ and $q_{2}$ is defined as:
$\theta=2\cdot\arccos\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left|q_{1}\cdot q_{2}}\right|}\right),$ (13)
where $q_{1}$ and $q_{2}$ are normalized quaternions, and $\cdot$ denotes the quaternion dot product. The Euclidean distance $d$ between two translation vectors $\mathbf{t}_{1}=(x_{1},y_{1},z_{1})$ and $\mathbf{t}_{2}=(x_{2},y_{2},z_{2})$ is given by:
$d=\sqrt{(x_{2}-x_{1})^{2}+(y_{2}-y_{1})^{2}+(z_{2}-z_{1})^{2}}.$ (14)
Appendix F Real-World Experiment Details
Our real-world hardware setup is based on the open-source Low-Cost-Robot project (Koch, 2024). We utilize two Intel RealSense D415 cameras for image capture. A visualization of our platform is provided in Fig. 8. For teleoperation, policy training, and evaluation, we leverage the open-source RealRobot project (Contributors, 2024). The policy used is the ACT policy (Zhao et al., 2023).
For each task, we collect 50 demonstrations, and during evaluation, we conduct 25 rollouts, each with randomized object locations and orientations. The model is trained for 10,000 epochs using four NVIDIA A100 GPUs. We employ the AdamW optimizer with a learning rate of $5\times 10^{-5}$ and a weight decay of 0.05. Additionally, a OneCycle learning rate scheduler is used, with a starting percentage of 0.1, a division factor of 10, and a final division factor of 100.
Appendix G Additional Reinforcement Learning Experiments
We conduct additional RL experiments following the settings in Hu et al. (2023) to use DrQ-v2 (Yarats et al., 2021), a state-of-the-art off-policy actor-critic approach for continuous vision-based control. We train some RL experiments with different pre-trained vision representations with ViT-Base architectures. The vision encoders are frozen during RL training. Five tasks in the Meta-World benchmark are chosen, as shown below. We train for a total of 1.1M frames and all other hyper-parameters including random seeds are kept as default and same. We run three seeds for each experiment. We report the evaluation success rate and episode reward below in Tab. 9. The reward curves are visualized in Fig. 7. From the results, it is evident that the reinforcement learning outcomes exhibit high variance. Nevertheless, overall, our 3D spatial-aware representation outperforms other representation learning methods.
Table 9: Reinforcement learning comparison results.
Meta-World RL Task Method (ViT-B) Success Rate Episode Reward
button-press-topdown-v2 CLIP 0.93 653.97
DINOv2 1.00 746.04
MAE 0.46 517.54
MoCoV3 0.99 749.93
SPA (Ours) 1.00 778.47
hammer-v2 CLIP 0.00 401.41
DINOv2 0.67 746.74
MAE 0.66 720.19
MoCoV3 0.59 645.46
SPA (Ours) 1.00 870.32
lever-pull-v2 CLIP 0.00 478.18
DINOv2 0.00 694.73
MAE 0.00 540.44
MoCoV3 0.23 598.54
SPA (Ours) 0.15 646.33
coffee-pull-v2 CLIP 0.00 181.40
DINOv2 0.00 180.72
MAE 0.00 184.56
MoCoV3 0.00 225.73
SPA (Ours) 0.00 262.11
drawer-close-v2 CLIP 1.00 1228.90
DINOv2 1.00 1236.30
MAE 1.00 1233.91
MoCoV3 1.00 1233.46
SPA (Ours) 1.00 1235.81
Mean CLIP 0.39 588.77
DINOv2 0.53 720.91
MAE 0.42 639.33
MoCoV3 0.56 690.63
SPA (Ours) 0.63 758.61
Figure 7: Reinforcement learning reward curves visualization.
Appendix H Additional Monocular Grasp Pose Detection Experiments
We conduct a monocular grasp pose detection experiment to further investigate more complex robotics learning paradigms. We follow similar settings in Gou et al. (2021), which train a neural network to detect the 7-DoF grasp poses on monocular image observations. The experiment is conducted on GraspNet-1Billion (Fang et al., 2020), a large-scale real-world object grasping benchmark. We follow the hyper-parameters and setups in the official implementation, except that we replace the default ResNet with different pre-trained ViT models for feature extraction. All pre-trained representations are with ViT-Base architecture and are frozen during training. We report the overall Top-K accuracy on the test set below. The results align well with our findings and indicate that SPA also outperforms other representation learning methods in the monocular grasp pose detection task.
Table 10: Overall top-k monocular grasp pose detection accuracy of various methods (ViT-Base).
Method (ViT-Base) CLIP DINOv2 MoCoV3 MAE SPA
Overall Accuracy 21.10 22.08 29.39 31.03 31.20
Appendix I Additional Ablation Study on Neural Rendering
To clarify the contribution of neural rendering to the overall performance of SPA, we conducted an additional ablation study. In this study, we maintained all settings identical—data loading, training techniques, hyperparameters, and the encoder—while replacing the volume neural rendering decoder with a multiview transformer-based decoder, similar to the MAE decoder. This alternative decoder receives masked patches filled with mask tokens corresponding to multiview images. Additional camera pose embeddings are added, and attention layers are used to fuse the multiview information and reconstruct RGB and depth images. We refer to this baseline as MV-MAE. It was trained on the ScanNet dataset without semantic supervision, ensuring a fair comparison with the result in the last line of Tab. 7. The results from this experiment demonstrate that neural rendering is crucial for incorporating explicit 3D spatial information. Simple multiview attention-based interaction, as used in MV-MAE, does not perform as effectively in learning 3D spatial awareness.
Table 11: Additional Ablation Study on Neural Rendering. The models are evaluated on two subsets of the VC-1 benchmark. The model architectures are both ViT-base.
Method Meta-World DMControl
MV-MAE 84.8$\pm$5.8 59.6$\pm$3.2
SPA 88.0$\pm$4.5 61.5$\pm$3.4
Figure 8: Real-world hardware platform.
Table 12: All results on Vc-1 benchmarks.
Benchmark AD MW DMC TF
Methods Seed Relo- cate Pen Button Press Topdown Drawer Open Bin Picking Ham- mer Assem- bly Walker Stand Walker Walk Reacher Easy Cheetah Run Finger Spin Reach Cube Move Cube
ViT-L Methods
MoCoV3 100 40.00 92.00 88.00 100.00 88.00 100.00 88.00 84.88 57.59 92.29 56.28 70.49 84.37 61.20
200 36.00 80.00 88.00 100.00 80.00 100.00 84.00 82.95 55.02 92.08 43.17 69.49 84.20 61.26
300 28.00 76.00 84.00 100.00 68.00 92.00 72.00 81.42 53.59 91.96 41.27 68.23 84.09 64.24
MAE 100 36.00 84.00 84.00 100.00 88.00 100.00 100.00 951.27 680.69 976.50 482.47 703.30 85.46 59.46
200 36.00 80.00 84.00 100.00 76.00 96.00 96.00 933.67 676.92 952.20 49.22 695.00 86.88 59.45
300 32.00 80.00 68.00 100.00 72.00 98.00 88.00 873.53 659.41 895.60 501.91 691.80 85.26 61.69
DINOV2 100 32.00 68.00 68.00 100.00 84.00 100.00 88.00 87.01 56.52 94.50 26.98 70.87 86.16 50.84
200 28.00 68.00 60.00 100.00 80.00 96.00 80.00 86.00 53.97 89.97 21.84 68.78 86.87 50.78
300 28.00 60.00 60.00 100.00 80.00 92.00 72.00 82.41 51.69 88.36 21.34 67.41 86.05 50.17
CLIP 100 24.00 80.00 28.00 100.00 88.00 100.00 84.00 66.16 43.94 90.71 18.40 68.30 73.28 41.09
200 24.00 72.00 24.00 100.00 84.00 100.00 80.00 64.04 34.51 88.26 16.52 66.68 75.11 33.45
300 24.00 68.00 16.00 100.00 84.00 96.00 72.00 52.70 31.60 85.17 14.51 67.49 74.73 38.95
EVA 100 44.00 84.00 72.00 100.00 76.00 100.00 100.00 77.92 51.64 98.17 31.04 70.17 82.56 52.04
200 40.00 76.00 84.00 100.00 72.00 100.00 100.00 77.63 50.81 86.66 19.37 67.43 81.72 52.07
300 32.00 72.00 96.00 100.00 68.00 96.00 96.00 77.21 47.71 88.41 29.19 67.43 82.13 52.70
InternViT- 300M 100 40.00 72.00 80.00 100.00 72.00 100.00 84.00 70.04 30.44 80.80 16.70 67.05 78.59 53.67
200 28.00 72.00 68.00 100.00 76.00 96.00 84.00 67.63 31.55 82.33 19.39 67.57 77.07 55.21
300 28.00 80.00 60.00 100.00 72.00 96.00 72.00 66.95 29.28 81.87 18.80 68.55 78.27 48.58
InternViT- 6B 100 32.00 72.00 88.00 100.00 80.00 100.00 84.00 88.53 70.02 93.09 26.54 70.62 85.96 57.52
200 40.00 76.00 84.00 100.00 76.00 100.00 80.00 85.28 60.09 90.86 22.84 69.04 86.30 54.11
300 60.00 80.00 80.00 100.00 88.00 100.00 76.00 81.88 59.17 87.87 21.53 67.20 85.68 54.86
MVP 100 32.00 84.00 96.00 100.00 96.00 100.00 100.00 84.88 57.59 92.29 56.28 70.49 84.37 61.20
200 28.00 76.00 92.00 100.00 84.00 100.00 96.00 82.95 55.02 92.08 43.17 69.49 84.20 61.26
300 24.00 76.00 84.00 100.00 68.00 100.00 88.00 81.42 53.59 91.96 41.27 68.23 84.09 64.24
VC-1 100 32.00 84.00 84.00 100.00 76.00 96.00 96.00 82.36 55.33 98.09 35.31 72.60 83.36 58.00
200 28.00 80.00 68.00 100.00 72.00 92.00 84.00 80.21 53.90 89.83 34.10 70.15 83.17 61.00
300 24.00 76.00 76.00 100.00 96.00 88.00 84.00 68.62 50.13 87.89 31.18 70.11 82.75 57.16
SPA-L 100 40.00 88.00 76.00 100.00 92.00 100.00 100.00 94.19 66.34 95.57 52.53 73.95 87.37 56.68
200 44.00 76.00 84.00 100.00 88.00 100.00 84.00 92.28 60.60 81.43 44.99 71.83 87.26 64.35
300 36.00 76.00 96.00 100.00 88.00 96.00 96.00 87.87 51.75 83.86 39.10 70.91 87.62 58.02
ViT-B Methods and Others
STP-B 100 20.00 80.00 88.00 100.00 84.00 100.00 96.00 77.02 45.34 87.97 40.01 72.72 80.41 54.66
200 28.00 76.00 92.00 100.00 84.00 100.00 80.00 71.50 33.60 84.08 34.30 72.18 80.13 57.97
300 32.00 76.00 88.00 100.00 72.00 100.00 96.00 71.44 42.86 79.67 39.17 69.12 80.65 53.95
R3M-B 100 20.00 92.00 52.00 96.00 32.00 88.00 48.00 668.49 301.54 842.90 256.56 678.00 75.08 45.66
200 12.00 76.00 48.00 96.00 32.00 88.00 44.00 634.62 256.82 661.40 198.63 660.90 75..62 48.09
300 12.00 76.00 32.00 88.00 28.00 76.00 40.00 633.39 211.90 585.50 188.15 657.30 74.54 45.59
Theia-B 100 32.00 76.00 88.00 100.00 80.00 96.00 96.00 72.90 43.97 82.09 37.02 70.50 84.55 55.62
200 36.00 80.00 60.00 100.00 84.00 100.00 84.00 79.05 56.99 94.36 39.59 70.22 83.27 54.59
300 24.00 72.00 80.00 100.00 72.00 96.00 100.00 79.64 54.39 82.89 39.09 72.00 84.01 54.43
Voltron-B 100 16.00 72.00 76.00 100.00 64.00 100.00 96.00 74.31 42.05 68.88 36.94 70.91 86.28 65.11
200 32.00 72.00 76.00 100.00 60.00 96.00 88.00 71.57 38.17 67.53 31.01 70.17 86.61 62.39
300 20.00 68.00 72.00 100.00 52.00 96.00 84.00 71.25 36.50 66.14 30.11 69.88 86.16 59.02
MAE-B 100 24.00 88.00 88.00 100.00 84.00 96.00 96.00 88.28 42.55 95.18 44.08 69.26 85.63 55.68
200 28.00 76.00 84.00 100.00 84.00 88.00 88.00 77.13 38.49 88.22 32.75 69.02 85.14 56.81
300 28.00 72.00 76.00 100.00 80.00 88.00 80.00 75.60 36.93 78.35 31.03 69.01 84.11 57.30
DINOV2-B 100 8.00 60.00 40.00 100.00 44.00 96.00 20.00 45.95 16.61 63.57 13.38 60.11 74.07 36.29
200 8.00 68.00 40.00 100.00 64.00 88.00 16.00 37.96 15.81 51.44 12.59 59.56 74.18 32.14
300 12.00 64.00 48.00 100.00 64.00 88.00 4.00 32.43 14.31 36.01 11.67 56.54 73.77 36.53
VC-1-B 100 20.00 76.00 76.00 100.00 76.00 100.00 76.00 72.35 43.14 92.77 27.31 68.67 84.19 62.00
200 32.00 80.00 68.00 100.00 76.00 100.00 92.00 81.83 44.05 83.62 27.80 70.98 83.88 59.63
300 24.00 68.00 80.00 100.00 80.00 88.00 88.00 83.01 41.25 77.60 28.53 70.89 84.76 59.51
RADIO 100 28.00 76.00 48.00 100.00 72.00 100.00 84.00 87.84 62.72 96.53 15.71 67.88 85.70 57.52
200 36.00 76.00 44.00 100.00 72.00 96.00 52.00 80.26 57.39 95.93 15.26 67.51 85.67 57.81
300 44.00 72.00 32.00 100.00 40.00 92.00 48.00 79.62 53.51 89.16 14.80 66.57 85.64 58.14
E-RADIO 100 32.00 84.00 64.00 100.00 84.00 96.00 96.00 71.47 53.41 93.01 50.19 70.75 87.17 46.97
200 32.00 84.00 60.00 100.00 68.00 88.00 96.00 68.80 44.56 88.54 33.19 70.46 87.39 50.72
300 28.00 80.00 60.00 100.00 72.00 88.00 80.00 65.96 33.56 98.14 32.64 69.18 87.09 51.31
SPA-B 100 20.00 84.00 84.00 100.00 88.00 100.00 100.00 80.50 45.08 91.38 48.90 71.16 86.04 59.03
200 28.00 80.00 68.00 100.00 84.00 100.00 84.00 79.71 46.65 85.75 40.84 71.01 86.16 60.05
300 24.00 80.00 88.00 100.00 92.00 100.00 92.00 74.70 48.97 81.60 34.92 71.16 85.16 61.94
Table 13: All results on Franka Kitchen.
Task View Seed MoCoV3 MAE DINOV2 CLIP EVA InternViT- 300M MVP VC-1 SPA
Task 1 Left 100 86.00 76.00 84.00 72.00 78.00 74.00 66.00 74.00 84.00
200 78.00 78.00 74.00 72.00 76.00 72.00 58.00 74.00 92.00
300 80.00 80.00 78.00 62.00 82.00 70.00 64.00 74.00 80.00
Right 100 82.00 80.00 86.00 78.00 78.00 72.00 82.00 78.00 86.00
200 88.00 62.00 90.00 70.00 86.00 86.00 86.00 84.00 72.00
300 86.00 82.00 92.00 82.00 86.00 76.00 92.00 78.00 86.00
Task 2 Left 100 60.00 56.00 48.00 26.00 40.00 22.00 40.00 32.00 48.00
200 64.00 60.00 46.00 44.00 40.00 32.00 32.00 42.00 60.00
300 58.00 54.00 40.00 26.00 32.00 34.00 30.00 50.00 66.00
Right 100 62.00 54.00 56.00 26.00 44.00 26.00 32.00 54.00 48.00
200 64.00 54.00 60.00 36.00 40.00 24.00 28.00 56.00 42.00
300 64.00 52.00 50.00 38.00 40.00 30.00 34.00 44.00 42.00
Task 3 Left 100 16.00 24.00 18.00 18.00 22.00 24.00 6.00 24.00 28.00
200 28.00 20.00 14.00 18.00 20.00 16.00 6.00 30.00 38.00
300 22.00 16.00 14.00 10.00 26.00 22.00 8.00 26.00 30.00
Right 100 46.00 26.00 38.00 22.00 14.00 8.00 32.00 12.00 10.00
200 48.00 22.00 38.00 24.00 18.00 4.00 32.00 12.00 12.00
300 54.00 34.00 52.00 14.00 12.00 6.00 26.00 14.00 16.00
Task 4 Left 100 32.00 36.00 26.00 22.00 34.00 12.00 16.00 36.00 22.00
200 30.00 30.00 32.00 14.00 20.00 8.00 14.00 24.00 10.00
300 24.00 46.00 28.00 14.00 32.00 4.00 20.00 36.00 16.00
Right 100 38.00 24.00 28.00 22.00 32.00 12.00 26.00 12.00 30.00
200 42.00 24.00 24.00 24.00 24.00 12.00 30.00 8.00 38.00
300 46.00 16.00 32.00 28.00 36.00 16.00 26.00 12.00 30.00
Task 5 Left 100 36.00 18.00 8.00 16.00 24.00 22.00 26.00 28.00 20.00
200 30.00 24.00 8.00 10.00 24.00 16.00 20.00 22.00 16.00
300 22.00 22.00 10.00 12.00 16.00 14.00 28.00 30.00 18.00
Right 100 24.00 46.00 20.00 4.00 14.00 10.00 26.00 22.00 30.00
200 24.00 30.00 16.00 8.00 18.00 18.00 22.00 22.00 26.00
300 14.00 36.00 16.00 12.00 10.00 12.00 20.00 16.00 22.00
Table 14: All results on Meta-World.
Method MoCoV3 MAE DINOV2 CLIP EVA InternViT- 300M MVP VC-1 SPA
Seed 100 200 300 100 200 300 100 200 300 100 200 300 100 200 300 100 200 300 100 200 300 100 200 300 100 200 300
ButtonPressWall 100 100 100 100 100 100 95 90 95 100 100 100 95 95 100 95 100 95 100 100 100 100 100 100 100 100 100
DoorClose 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100
DoorUnlock 70 65 80 70 65 85 35 35 30 60 50 60 85 85 90 80 65 75 75 70 85 80 75 90 80 75 80
DrawerClose 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100
DrawerOpen 80 70 85 85 65 75 60 40 60 90 80 80 60 55 75 70 75 75 95 75 85 85 85 80 90 60 75
FaucetClose 70 80 55 60 80 60 55 65 35 70 80 50 65 75 60 50 65 50 70 75 60 65 75 100 70 80 65
PlateSlide 90 95 100 95 100 100 65 70 80 85 95 80 100 100 100 85 95 90 100 100 100 100 95 100 95 95 95
PlateSlideBack 80 65 85 85 65 85 90 75 90 85 75 90 80 65 90 85 80 90 85 70 90 80 70 90 80 70 85
PlateSlideSide 85 90 95 95 90 95 90 85 85 80 85 95 100 95 100 90 95 90 80 90 95 100 95 90 90 90 95
WindowClose 100 100 100 100 100 100 70 90 100 100 100 95 95 100 100 100 100 100 100 100 100 100 100 100 100 100 100
Basketball 85 95 95 95 100 100 70 55 65 85 85 70 95 85 80 90 80 95 95 100 95 90 100 95 95 100 100
BinPicking 30 45 40 20 10 35 10 15 10 30 30 25 20 5 45 10 10 10 25 20 15 30 30 30 40 25 30
BoxClose 80 80 80 75 80 70 35 45 30 80 70 60 55 70 55 60 65 40 80 80 65 80 95 60 80 80 65
CoffeePush 45 50 40 40 55 30 30 25 15 45 40 55 45 35 40 45 30 25 25 45 25 30 45 55 35 40 30
Assembly 70 60 55 55 65 45 30 35 25 45 55 50 45 50 40 30 25 30 60 60 45 60 60 50 50 55 50
Disassemble 40 55 50 30 45 45 30 20 45 55 50 60 30 50 50 40 45 50 40 45 45 40 30 35 40 45 55
PushWall 25 35 30 20 30 40 40 30 45 35 30 35 25 35 30 15 15 25 30 35 35 55 55 60 30 40 45
ShelfPlace 35 35 20 25 45 20 30 30 35 30 35 30 25 20 15 15 15 15 25 25 15 25 15 15 15 35 15
DoorOpen 95 90 90 100 95 95 95 80 95 85 75 90 80 90 100 50 55 60 80 75 95 95 95 95 85 100 95
ButtonPress 75 85 85 85 100 100 80 95 85 55 70 75 80 90 90 85 90 80 85 90 95 80 90 85 100 95 100
SweepInto 45 45 40 55 55 45 50 50 40 45 45 40 45 25 30 35 25 30 45 50 40 55 50 50 50 50 45
DoorLock 100 85 85 90 100 100 95 90 85 85 85 75 95 100 95 80 90 95 85 100 95 100 100 90 80 95 85
ReachWall 70 70 80 75 85 70 85 80 85 90 90 80 65 75 85 60 55 55 75 65 75 75 80 80 75 80 70
Hammer 25 45 30 30 30 30 40 45 35 25 35 25 30 35 20 20 20 20 30 35 30 45 40 55 30 40 30
StickPush 95 95 100 80 90 95 90 90 90 75 85 85 90 85 100 60 80 85 90 95 90 85 85 95 90 90 85
ButtonPressTopdown 80 80 75 80 85 80 80 90 90 80 90 70 55 65 55 45 55 55 80 85 80 75 80 70 80 85 80
HandlePressSide 100 100 100 100 100 100 95 100 90 80 100 90 100 100 95 85 100 90 95 100 100 75 100 80 90 100 100
PlateSlideBackSide 100 100 100 100 95 95 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100 100
Sweep 50 80 70 35 60 60 65 85 95 60 85 75 35 50 55 15 60 35 35 70 65 35 65 65 30 65 55
ButtonPressTopdownWall 45 70 80 45 75 75 70 75 75 45 60 70 30 45 65 20 55 45 30 60 70 50 85 80 45 65 75
HandlePress 85 95 95 80 100 75 75 100 80 90 100 90 85 100 85 65 90 75 80 95 75 85 100 90 85 100 80
Push 25 30 30 25 30 30 30 25 40 25 15 30 30 25 20 25 20 20 25 20 25 40 30 25 30 15 35
CoffeePull 55 55 55 40 45 40 20 30 20 50 70 40 40 45 45 40 40 25 55 55 40 55 55 60 55 55 55
DialTurn 80 65 80 85 75 75 40 30 35 80 95 90 65 65 80 70 70 55 70 65 75 80 95 75 85 85 75
Reach 90 75 80 90 75 80 70 75 85 95 95 100 85 80 75 95 80 90 80 70 75 70 80 80 85 70 85
CoffeeButton 85 95 75 100 100 95 85 80 60 95 100 85 90 100 95 90 85 85 100 100 90 90 100 80 100 100 100
PickPlaceWall 45 35 65 40 25 45 15 10 20 35 40 50 20 25 35 30 25 25 25 35 40 35 20 45 40 35 55
StickPull 35 35 25 15 40 20 25 10 5 45 40 45 25 35 15 25 25 15 15 30 25 30 30 30 25 35 30
HandInsert 35 30 30 30 25 25 20 20 20 45 45 40 20 25 20 25 30 25 40 40 35 40 30 40 40 50 40
PegInsertSide 40 35 40 50 35 45 25 15 10 45 30 20 45 25 30 30 20 25 45 45 30 50 25 35 55 45 60
PickPlace 35 30 30 25 45 15 15 10 10 25 15 30 25 30 25 30 10 30 20 35 25 25 20 20 25 30 40
FaucetOpen 95 95 100 100 100 100 80 80 100 100 95 100 95 95 100 80 85 95 100 100 100 100 100 100 95 100 100
PushBack 65 70 60 40 55 40 15 15 25 30 45 25 35 35 45 15 35 25 40 45 45 45 55 25 35 55 50
LeverPull 70 80 80 80 70 75 15 30 35 55 85 80 70 65 70 60 55 55 65 80 65 65 80 70 85 70 80
HandlePull 85 85 80 85 80 85 45 55 40 75 75 80 80 60 65 45 70 65 80 90 80 70 85 75 100 90 90
Soccer 25 40 35 50 30 25 15 10 15 45 40 30 35 35 25 20 20 30 30 30 35 20 20 20 25 50 25
WindowOpen 65 80 80 55 80 85 60 50 65 50 65 60 65 80 75 55 75 75 60 70 70 60 70 75 55 85 65
PickOutOfHole 65 75 80 65 65 60 60 55 60 70 70 50 60 55 50 60 55 55 65 55 60 70 75 60 65 60 75
Table 15: All results on RLBench.
Method MoCoV3 MAE DINOV2 CLIP EVA InternViT 300M MVP VC-1 SPA
Group 1
basketball in hoop 100 100 100 100 100 100 100 100 100
put rubbish in bin 100 100 96 96 96 100 96 100 100
meat off grill 100 100 100 100 100 100 100 100 100
meat on grill 80 76 76 68 80 72 68 76 80
slide block to target 0 84 96 24 4 0 100 100 4
reach and drag 100 96 88 100 96 100 96 100 100
take frame off hanger 88 88 92 88 84 84 88 88 96
water plants 64 60 28 64 60 44 52 60 68
hang frame on hanger 8 4 0 4 8 8 12 4 4
wipe desk 0 0 0 0 0 0 0 0 0
stack blocks 60 72 72 68 56 60 84 68 68
reach target 60 96 88 100 96 80 92 96 92
push button 100 100 100 100 100 100 100 100 100
lamp on 88 68 84 88 52 80 28 88 64
toilet seat down 100 100 100 100 100 100 96 96 100
close laptop lid 96 96 96 96 84 80 80 96 100
open box 12 12 20 4 16 4 0 12 16
open drawer 88 96 92 100 88 88 92 96 96
pick up cup 92 92 88 96 96 88 96 96 96
turn tap 88 84 84 96 88 92 96 100 100
take usb out of computer 100 100 100 100 100 100 100 88 100
play jenga 96 96 96 100 96 100 96 96 96
insert onto square peg 28 84 80 44 88 40 64 92 84
take umbrella out of umbrella stand 92 100 100 92 100 96 100 100 100
insert usb in computer 12 20 20 24 24 20 16 8 68
straighten rope 56 44 72 80 48 72 52 60 84
turn oven on 96 96 96 96 96 96 100 100 100
change clock 64 68 48 68 64 72 64 60 68
close microwave 100 100 100 100 100 100 100 100 100
close fridge 80 92 92 88 92 96 88 92 100
close grill 96 96 96 96 96 96 100 100 96
open grill 100 100 100 100 100 100 96 100 100
unplug charger 44 32 48 36 48 40 40 44 44
press switch 92 92 88 72 76 84 76 88 92
take money out safe 100 96 100 100 100 100 100 100 100
Group 2
change channel 0 8 4 0 0 4 0 0 4
tv on 4 8 0 4 4 8 4 4 8
push buttons 12 4 4 0 0 0 0 12 4
stack wine 12 16 40 4 12 0 28 8 28
scoop with spatula 0 0 0 0 0 0 0 0 0
place hanger on rack 0 0 0 0 0 0 0 0 0
move hanger 0 0 0 0 0 0 0 0 0
sweep to dustpan 92 96 96 96 92 100 100 88 96
take plate off colored dish rack 96 100 96 92 84 96 88 92 96
screw nail 52 36 36 36 36 52 32 32 48
take shoes out of box 20 28 24 36 40 12 32 36 36
slide cabinet open and place cups 0 0 0 0 0 4 0 0 4
lamp off 100 96 96 100 96 96 100 100 100
pick and lift 88 96 92 96 92 80 96 96 96
take lid off saucepan 100 100 100 100 100 100 100 100 100
close drawer 100 100 100 100 96 100 100 100 100
close box 92 92 96 96 100 96 100 96 100
phone on base 100 100 100 100 100 96 100 100 100
toilet seat up 80 88 100 88 88 80 88 92 96
put books on bookshelf 12 24 24 28 28 20 20 28 16
beat the buzz 88 92 96 88 92 84 88 88 100
stack cups 40 56 52 52 48 56 64 68 64
put knife on chopping board 72 76 68 72 80 88 80 76 80
place shape in shape sorter 20 36 32 28 36 20 44 36 56
take toilet roll off stand 100 92 76 96 92 88 84 92 96
put umbrella in umbrella stand 8 0 12 12 0 4 12 8 12
setup checkers 76 80 68 68 88 92 92 80 80
open window 96 96 100 100 96 100 96 100 100
open wine bottle 80 100 88 92 92 88 96 88 88
open microwave 100 100 88 96 100 80 96 100 100
put money in safe 96 100 88 92 100 96 100 100 100
open door 100 96 96 96 96 96 84 96 96
close door 32 68 56 60 80 20 24 20 60
open fridge 44 52 48 44 36 64 52 32 64
open oven 8 4 12 8 4 20 4 4 16
plug charger in power supply 32 36 32 24 44 36 24 32 60
Table 16: All results on LIBERO-OBJECT, LIBERO-SPATIAL, LIBERO-GOAL, LIBERO-10.
MoCoV3 MAE DINOV2 CLIP EVA InternViT- 300M InternViT- 6B MVP VC-1 SPA
Seed 100 200 300 100 200 300 100 200 300 100 200 300 100 200 300 100 200 300 100 200 300 100 200 300 100 200 300 100 200 300
LIBERO-OBJECT
0 0.65 0.60 0.65 0.65 0.45 0.55 0.65 0.80 0.85 0.80 0.75 0.65 1.00 0.70 0.95 0.80 0.65 0.60 0.70 0.85 0.50 0.80 0.90 0.65 0.80 0.50 0.60 0.90 0.95 0.95
1 0.35 0.35 0.55 0.90 0.75 0.80 0.30 0.50 0.75 0.40 0.30 0.05 0.65 0.30 0.70 0.15 0.40 0.20 0.60 0.25 0.45 0.05 0.80 0.60 0.40 0.65 0.45 0.65 0.70 0.45
2 0.90 0.85 0.95 0.90 0.40 0.95 0.85 0.50 0.90 0.70 0.80 0.75 0.85 0.75 0.75 0.90 0.85 0.80 0.85 0.45 0.85 0.80 0.85 0.90 1.00 0.95 0.95 0.90 0.95 0.80
3 0.55 0.70 0.65 0.90 0.15 0.90 0.30 0.65 0.90 0.25 0.45 0.60 0.80 0.80 0.90 0.75 0.70 0.40 1.00 0.50 0.55 0.70 0.65 0.85 0.95 0.75 0.60 0.70 0.90 0.90
4 0.65 0.85 0.85 0.80 0.90 0.75 0.75 0.55 0.75 0.35 0.75 0.65 0.95 0.75 1.00 0.90 1.00 0.85 0.90 0.70 0.80 0.80 0.75 0.70 0.90 0.85 0.90 0.90 1.00 0.95
5 0.50 0.70 0.80 0.70 0.35 0.60 0.55 0.75 0.60 0.25 0.70 0.45 0.75 0.75 0.65 0.85 0.60 0.75 0.60 0.35 0.50 0.55 0.40 0.80 0.65 0.70 0.70 0.25 0.15 0.65
6 0.35 0.50 0.65 0.60 0.65 0.65 0.55 0.70 0.70 0.35 0.55 0.60 0.40 0.35 0.25 0.65 0.60 0.55 0.30 0.10 0.35 0.25 0.50 0.65 0.50 0.50 0.30 0.50 0.70 0.80
7 0.75 0.75 0.80 0.90 0.40 0.75 0.55 0.30 0.70 0.40 0.35 0.40 0.55 0.75 0.70 0.80 0.40 0.60 0.60 0.65 0.70 0.60 0.45 0.65 0.80 0.75 0.50 0.80 0.75 0.65
8 0.50 0.95 0.90 1.00 0.95 1.00 0.50 0.35 0.50 0.45 0.45 0.35 1.00 0.75 0.85 0.70 0.65 0.75 0.50 0.40 0.75 0.65 0.55 0.70 0.80 0.90 0.50 0.85 0.95 0.90
9 0.45 0.50 0.40 0.60 0.65 0.95 0.80 0.90 0.95 0.50 0.70 0.30 0.85 0.75 0.75 0.85 0.95 0.65 0.90 0.60 0.15 0.65 0.60 0.30 0.70 0.70 0.65 0.60 0.95 0.90
LIBERO-SPATIAL
0 0.35 0.55 0.45 0.45 0.40 0.70 0.65 0.50 0.60 0.25 0.20 0.35 0.70 0.75 0.65 0.55 0.65 0.55 0.55 0.50 0.30 0.75 0.75 0.60 0.35 0.55 0.60 0.45 0.50 0.35
1 0.65 0.70 0.70 0.80 0.80 0.50 0.55 0.30 0.35 0.75 0.75 0.70 0.55 0.70 0.25 0.35 0.50 0.50 1.00 1.00 0.90 0.60 0.40 0.60 0.45 0.65 0.80 0.65 0.65 0.85
2 0.55 0.50 0.50 0.35 0.60 0.40 0.20 0.05 0.55 0.10 0.00 0.40 0.70 0.80 0.50 0.70 0.75 0.60 0.75 0.60 0.20 0.85 0.55 0.75 0.45 0.45 0.70 0.50 0.50 0.40
3 0.50 0.70 0.75 0.55 0.60 0.75 0.80 0.70 0.95 0.15 0.40 0.30 0.85 0.90 0.85 0.35 0.50 0.40 0.40 0.30 0.15 0.95 0.55 0.60 0.50 0.70 0.65 0.55 0.85 0.60
4 0.15 0.15 0.20 0.55 0.70 0.80 0.50 0.05 0.45 0.35 0.30 0.20 0.45 0.55 0.40 0.35 0.25 0.40 0.25 0.15 0.15 0.60 0.50 0.70 0.60 0.60 0.80 0.70 0.70 0.50
5 0.45 0.10 0.10 0.65 0.40 0.30 0.30 0.20 0.35 0.55 0.45 0.45 0.65 0.50 0.45 0.40 0.30 0.70 0.55 0.60 0.60 0.55 0.30 0.25 0.05 0.05 0.15 0.35 0.35 0.30
6 0.30 0.35 0.45 0.55 0.25 0.95 0.40 0.30 0.40 0.20 0.25 0.10 0.75 0.70 0.85 0.45 0.55 0.55 0.40 0.35 0.05 0.45 0.75 0.65 0.60 0.70 0.35 0.35 0.45 0.30
7 0.10 0.20 0.25 0.50 0.35 0.45 0.05 0.00 0.10 0.30 0.15 0.25 0.30 0.30 0.20 0.30 0.60 0.65 0.15 0.05 0.05 0.60 0.65 0.60 0.10 0.20 0.40 0.40 0.15 0.45
8 0.55 0.70 0.50 0.35 0.65 0.70 0.40 0.15 0.55 0.30 0.55 0.30 0.85 0.70 0.60 0.30 0.55 0.60 0.65 0.40 0.35 0.70 0.40 0.55 0.70 0.60 0.70 0.75 0.70 0.40
9 0.55 0.05 0.10 0.85 0.75 0.50 0.20 0.05 0.25 0.20 0.20 0.20 0.55 0.50 0.30 0.35 0.50 0.30 0.45 0.40 0.35 0.45 0.45 0.30 0.50 0.55 0.65 0.35 0.50 0.45
LIBERO-GOAL
0 0.45 0.70 0.75 0.70 0.85 0.80 0.15 0.10 0.30 0.25 0.40 0.35 0.70 0.60 0.60 0.75 0.65 0.75 0.25 0.35 0.35 0.75 0.60 0.95 0.45 0.85 1.00 0.85 1.00 0.85
1 0.70 0.60 0.80 0.65 0.50 0.90 0.25 0.55 0.25 0.20 0.15 0.25 0.70 0.80 0.80 0.90 0.90 1.00 0.40 0.15 0.15 0.90 0.80 0.95 0.65 0.65 0.65 1.00 0.85 0.90
2 0.50 0.20 0.15 0.10 0.40 0.35 0.10 0.05 0.15 0.30 0.25 0.30 0.65 0.75 0.75 0.40 0.75 0.45 0.50 0.35 0.35 0.45 0.25 0.65 0.40 0.60 0.35 0.50 0.55 0.35
3 0.75 0.45 0.60 0.40 0.75 0.55 0.20 0.10 0.10 0.05 0.20 0.55 0.30 0.15 0.15 0.30 0.50 0.65 0.20 0.25 0.25 0.70 0.70 0.15 0.75 0.55 0.50 0.65 0.35 0.80
4 0.20 0.25 0.05 0.35 0.40 0.25 0.10 0.00 0.05 0.40 0.30 0.15 0.15 0.10 0.10 0.20 0.15 0.10 0.15 0.20 0.20 0.55 0.60 0.25 0.15 0.30 0.30 0.30 0.35 0.35
5 0.10 0.75 0.80 0.60 0.85 0.80 0.65 0.50 0.50 0.35 0.45 0.50 0.80 0.75 0.75 0.70 0.55 0.45 0.55 0.45 0.45 0.80 0.75 0.85 0.65 0.75 0.80 0.80 0.65 0.65
6 0.45 0.05 0.15 0.00 0.10 0.05 0.00 0.05 0.00 0.10 0.10 0.00 0.00 0.65 0.65 0.50 0.40 0.30 0.00 0.00 0.00 0.15 0.35 0.70 0.25 0.50 0.45 0.40 0.30 0.35
7 0.25 0.75 0.90 0.80 0.65 1.00 0.45 0.45 0.35 0.50 0.65 0.80 1.00 1.00 1.00 1.00 1.00 0.95 0.70 0.85 0.85 0.95 1.00 0.95 1.00 0.95 0.70 0.95 1.00 1.00
8 0.50 0.80 0.75 0.85 0.55 0.90 0.45 0.40 0.20 0.60 0.25 0.50 0.90 0.35 0.35 0.95 0.65 0.55 0.65 0.25 0.25 0.70 0.65 0.70 0.50 0.75 0.55 0.80 0.65 0.80
9 0.10 0.65 0.60 0.50 0.20 0.50 0.10 0.00 0.10 0.10 0.10 0.00 0.15 0.70 0.70 0.60 0.40 0.20 0.15 0.35 0.35 0.30 0.50 0.55 0.20 0.35 0.70 0.55 0.60 0.45
LIBERO-10
0 0.15 0.20 0.10 0.15 0.25 0.10 0.00 0.05 0.10 0.00 0.05 0.05 0.25 0.35 0.10 0.35 0.10 0.25 0.15 0.15 0.00 0.05 0.15 0.20 0.10 0.45 0.25 0.05 0.10 0.05
1 0.25 0.20 0.20 0.30 0.15 0.25 0.15 0.15 0.15 0.40 0.30 0.15 0.65 0.10 0.60 0.15 0.50 0.45 0.00 0.25 0.35 0.15 0.10 0.15 0.20 0.40 0.15 0.25 0.05 0.45
2 0.70 0.60 0.75 0.30 0.60 0.75 0.55 0.45 0.50 0.25 0.45 0.40 0.75 0.55 0.65 0.45 0.80 0.55 0.70 0.80 0.75 0.75 0.65 0.55 0.85 1.00 0.90 0.70 0.80 0.50
3 0.50 0.80 0.55 0.55 0.60 0.80 0.40 0.45 0.45 0.60 0.65 0.60 0.80 0.90 0.75 0.75 0.65 0.50 0.75 0.60 0.60 0.75 0.70 0.65 0.80 0.70 0.70 0.70 0.90 0.70
4 0.25 0.20 0.05 0.35 0.25 0.30 0.10 0.10 0.05 0.20 0.05 0.05 0.15 0.10 0.15 0.15 0.15 0.05 0.15 0.20 0.15 0.25 0.20 0.35 0.30 0.25 0.30 0.40 0.30 0.25
5 0.40 0.60 0.75 0.55 0.70 0.80 0.50 0.65 0.75 0.40 0.40 0.30 0.85 0.65 0.75 0.75 0.55 0.75 0.45 0.55 0.45 0.60 0.70 0.75 0.80 0.90 0.60 0.70 0.70 0.45
6 0.20 0.25 0.10 0.40 0.35 0.40 0.05 0.20 0.05 0.25 0.15 0.15 0.20 0.30 0.35 0.10 0.15 0.20 0.20 0.25 0.05 0.40 0.30 0.35 0.30 0.10 0.20 0.20 0.20 0.15
7 0.40 0.30 0.25 0.50 0.50 0.50 0.10 0.30 0.60 0.30 0.40 0.20 0.45 0.30 0.25 0.40 0.35 0.35 0.35 0.70 0.25 0.35 0.30 0.25 0.30 0.25 0.30 0.50 0.45 0.40
8 0.10 0.10 0.15 0.10 0.30 0.20 0.35 0.20 0.05 0.10 0.10 0.10 0.15 0.20 0.10 0.20 0.00 0.25 0.05 0.20 0.05 0.10 0.30 0.20 0.25 0.30 0.25 0.25 0.05 0.15
9 0.20 0.60 0.35 0.40 0.65 0.30 0.35 0.25 0.45 0.45 0.45 0.30 0.40 0.70 0.50 0.50 0.45 0.60 0.50 0.25 0.40 0.40 0.55 0.50 0.00 0.00 0.00 0.50 0.65 0.50
Table 17: All results on LIBERO-90.
MoCoV3 MAE DINOV2 CLIP EVA InternViT- 300M InternViT- 6B MVP VC-1 SPA
Seed 100 200 300 100 200 300 100 200 300 100 200 300 100 200 300 100 200 300 100 200 300 100 200 300 100 200 300 100 200 300
LIBERO-90
0 0.95 0.85 0.90 1.00 0.90 0.80 0.80 1.00 0.60 0.90 0.80 0.80 1.00 1.00 1.00 0.90 0.80 0.85 0.75 0.80 0.95 0.95 0.95 1.00 0.95 1.00 0.95 1.00 1.00 0.95
1 0.60 0.35 0.60 0.35 0.50 0.15 0.50 0.55 0.30 0.40 0.65 0.35 0.70 0.50 0.25 0.30 0.45 0.40 0.25 0.35 0.55 0.80 0.55 0.30 0.40 0.50 0.05 0.65 0.40 0.50
2 0.85 0.50 0.80 0.55 0.55 0.20 0.65 0.60 0.30 0.45 0.30 0.50 0.35 0.50 0.70 0.85 0.65 0.80 0.25 0.35 0.30 0.45 0.70 0.70 0.75 0.55 0.35 0.70 0.85 0.60
3 0.10 0.10 0.00 0.05 0.00 0.00 0.05 0.00 0.00 0.15 0.00 0.00 0.00 0.10 0.15 0.00 0.05 0.05 0.00 0.00 0.05 0.10 0.00 0.05 0.05 0.10 0.00 0.10 0.10 0.00
4 0.40 0.05 0.20 0.30 0.25 0.30 0.15 0.40 0.55 0.40 0.40 0.35 0.10 0.25 0.15 0.40 0.05 0.25 0.20 0.45 0.40 0.30 0.40 0.15 0.25 0.05 0.15 0.15 0.15 0.35
5 0.05 0.05 0.05 0.10 0.05 0.20 0.00 0.20 0.00 0.05 0.05 0.20 0.25 0.25 0.10 0.10 0.05 0.30 0.20 0.10 0.25 0.05 0.15 0.10 0.10 0.05 0.05 0.35 0.00 0.15
6 0.10 0.00 0.00 0.00 0.00 0.05 0.05 0.05 0.10 0.05 0.10 0.00 0.00 0.00 0.05 0.00 0.00 0.00 0.05 0.10 0.10 0.00 0.05 0.00 0.00 0.00 0.00 0.00 0.05 0.05
7 0.35 0.30 0.65 0.20 0.60 0.30 0.35 0.25 0.40 0.50 0.60 0.35 0.60 0.10 0.20 0.15 0.25 0.10 0.40 0.25 0.45 0.50 0.20 0.40 0.20 0.30 0.25 0.30 0.30 0.65
8 0.10 0.15 0.00 0.05 0.20 0.10 0.15 0.25 0.10 0.20 0.10 0.10 0.10 0.05 0.00 0.05 0.00 0.10 0.20 0.15 0.20 0.10 0.00 0.20 0.05 0.15 0.20 0.05 0.05 0.15
9 0.30 0.25 0.35 0.50 0.25 0.30 0.35 0.60 0.70 0.25 0.20 0.50 0.25 0.10 0.50 0.10 0.10 0.25 0.60 0.25 0.30 0.25 0.15 0.45 0.25 0.05 0.35 0.25 0.20 0.25
10 0.50 0.75 0.50 0.50 0.60 0.55 0.65 0.60 0.60 0.90 0.45 0.55 0.40 0.85 0.35 0.35 0.05 0.25 0.45 0.45 0.65 0.40 0.50 0.55 0.45 0.75 0.40 0.40 0.35 0.35
11 0.45 0.35 0.75 0.45 0.70 0.65 0.35 0.20 0.15 0.40 0.70 0.55 0.80 0.25 0.70 0.50 0.50 0.10 0.35 0.25 0.45 0.80 0.60 0.95 0.70 0.75 0.60 0.60 0.60 0.65
12 0.15 0.15 0.10 0.15 0.15 0.05 0.20 0.20 0.15 0.10 0.05 0.05 0.10 0.25 0.05 0.05 0.00 0.00 0.25 0.30 0.10 0.15 0.10 0.10 0.20 0.25 0.10 0.05 0.10 0.15
13 0.20 0.35 0.30 0.15 0.30 0.20 0.30 0.35 0.10 0.30 0.40 0.35 0.30 0.10 0.45 0.20 0.35 0.40 0.25 0.15 0.55 0.30 0.30 0.15 0.45 0.10 0.10 0.10 0.20 0.10
14 0.05 0.10 0.00 0.30 0.30 0.20 0.10 0.10 0.15 0.15 0.40 0.20 0.25 0.35 0.10 0.15 0.15 0.05 0.20 0.15 0.10 0.20 0.35 0.10 0.20 0.10 0.20 0.15 0.15 0.10
15 0.60 0.75 0.45 0.70 0.50 0.65 0.35 0.50 0.55 0.45 0.65 0.40 0.70 0.75 0.40 0.40 0.65 0.50 0.35 0.55 0.45 0.70 0.60 0.55 0.80 0.80 0.70 0.65 0.80 0.55
16 0.05 0.20 0.00 0.30 0.15 0.05 0.10 0.10 0.05 0.10 0.00 0.10 0.20 0.20 0.15 0.15 0.15 0.20 0.05 0.00 0.10 0.15 0.05 0.10 0.00 0.15 0.15 0.05 0.10 0.15
17 0.05 0.15 0.15 0.10 0.25 0.05 0.05 0.10 0.05 0.05 0.00 0.05 0.05 0.20 0.15 0.10 0.10 0.15 0.00 0.10 0.00 0.20 0.10 0.20 0.15 0.10 0.00 0.25 0.10 0.10
18 0.45 0.40 0.60 0.40 0.75 0.65 0.30 0.35 0.40 0.45 0.25 0.35 0.25 0.35 0.60 0.40 0.05 0.70 0.60 0.50 0.35 0.35 0.25 0.45 0.30 0.60 0.35 0.60 0.35 0.55
19 0.30 0.30 0.25 0.35 0.40 0.20 0.20 0.05 0.35 0.45 0.45 0.30 0.30 0.35 0.25 0.15 0.25 0.20 0.35 0.30 0.15 0.55 0.30 0.40 0.40 0.45 0.35 0.45 0.20 0.35
20 0.85 0.75 0.80 1.00 1.00 0.95 0.95 1.00 1.00 0.75 0.85 0.30 1.00 1.00 1.00 0.95 0.95 0.90 1.00 0.50 0.80 0.90 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
21 0.40 0.20 0.40 0.35 0.25 0.30 0.25 0.40 0.20 0.25 0.10 0.45 0.30 0.70 0.05 0.00 0.05 0.10 0.35 0.10 0.30 0.40 0.15 0.30 0.30 0.70 0.60 0.65 0.40 0.60
22 0.90 0.95 0.95 1.00 0.85 0.95 0.25 0.60 0.40 0.75 0.75 0.75 0.95 1.00 0.95 0.85 0.95 0.60 0.45 0.25 0.25 0.90 1.00 1.00 0.90 0.90 0.95 1.00 0.95 1.00
23 0.15 0.05 0.15 0.05 0.10 0.00 0.05 0.10 0.05 0.00 0.00 0.05 0.00 0.00 0.00 0.00 0.00 0.00 0.20 0.25 0.05 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
24 0.80 0.30 0.85 0.85 0.50 0.80 0.60 0.50 0.65 0.70 0.45 0.60 0.70 0.70 0.80 0.40 0.65 0.60 0.55 0.80 0.45 0.65 0.60 0.90 0.90 0.80 0.80 0.90 0.80 0.75
25 1.00 0.80 0.85 1.00 1.00 0.90 0.75 0.90 0.90 0.80 0.95 0.90 0.90 1.00 0.95 0.70 0.70 0.85 0.95 0.60 0.65 1.00 0.85 1.00 1.00 1.00 0.90 1.00 1.00 1.00
26 0.15 0.20 0.25 0.25 0.40 0.40 0.05 0.30 0.40 0.45 0.05 0.15 0.05 0.30 0.15 0.25 0.40 0.20 0.25 0.15 0.20 0.20 0.30 0.60 0.25 0.45 0.25 0.25 0.20 0.20
27 0.30 0.15 0.20 0.35 0.35 0.10 0.05 0.10 0.00 0.35 0.05 0.10 0.05 0.20 0.05 0.05 0.15 0.00 0.10 0.10 0.05 0.35 0.20 0.30 0.10 0.45 0.40 0.10 0.40 0.20
28 0.90 0.90 1.00 0.95 0.70 0.70 0.80 0.50 0.80 0.90 0.75 0.90 0.95 1.00 0.85 0.90 0.85 1.00 0.50 0.60 0.45 0.85 0.75 0.90 0.75 0.95 0.60 0.90 0.65 0.90
29 0.15 0.50 0.35 0.60 0.55 0.20 0.50 0.50 0.40 0.50 0.50 0.30 0.65 0.30 0.30 0.15 0.30 0.40 0.25 0.35 0.25 0.35 0.50 0.25 0.60 0.60 0.10 0.30 0.35 0.70
30 0.15 0.25 0.15 0.60 0.35 0.35 0.35 0.10 0.50 0.25 0.20 0.45 0.30 0.70 0.20 0.10 0.15 0.20 0.50 0.20 0.25 0.00 0.05 0.20 0.10 0.25 0.10 0.40 0.25 0.40
31 0.70 0.60 0.80 0.70 0.75 0.60 0.45 0.70 0.75 0.95 0.65 0.95 0.80 0.75 0.45 0.50 0.55 0.35 0.95 0.80 0.85 0.80 0.70 0.75 0.75 0.75 0.45 0.70 0.90 0.80
32 0.30 0.05 0.20 0.05 0.10 0.05 0.00 0.00 0.05 0.35 0.10 0.10 0.20 0.10 0.15 0.10 0.15 0.10 0.05 0.15 0.05 0.20 0.25 0.10 0.05 0.10 0.05 0.20 0.10 0.05
33 0.50 0.55 0.40 0.15 0.30 0.30 0.10 0.20 0.35 0.30 0.25 0.30 0.00 0.50 0.40 0.35 0.20 0.25 0.25 0.25 0.30 0.20 0.35 0.40 0.35 0.45 0.65 0.15 0.15 0.15
34 0.30 0.35 0.30 0.40 0.40 0.25 0.35 0.40 0.15 0.40 0.40 0.50 0.10 0.40 0.10 0.15 0.05 0.25 0.35 0.30 0.50 0.25 0.30 0.30 0.55 0.25 0.05 0.05 0.30 0.10
35 0.65 0.40 0.60 0.85 0.95 0.75 0.80 0.80 0.60 0.65 0.55 0.90 0.90 1.00 0.80 0.10 0.20 0.55 0.85 0.75 0.70 0.85 0.80 0.85 0.25 0.80 0.55 1.00 0.70 1.00
36 0.05 0.10 0.15 0.05 0.00 0.00 0.00 0.25 0.05 0.05 0.05 0.00 0.15 0.20 0.10 0.15 0.10 0.25 0.00 0.00 0.10 0.20 0.00 0.15 0.30 0.05 0.20 0.05 0.20 0.10
37 0.35 0.30 0.40 0.55 0.20 0.25 0.65 0.50 0.35 0.30 0.60 0.60 0.70 0.80 0.60 0.45 0.55 0.45 0.50 0.55 0.55 0.45 0.45 0.50 0.45 0.75 0.35 0.75 0.50 0.55
38 0.50 0.30 0.45 0.55 0.50 0.35 0.25 0.15 0.30 0.50 0.45 0.30 0.50 0.20 0.35 0.35 0.55 0.45 0.35 0.50 0.35 0.70 0.40 0.55 0.70 0.65 0.30 0.70 0.45 0.45
39 0.80 0.80 0.75 0.45 0.70 0.60 0.60 0.55 0.65 0.60 0.65 0.70 0.65 0.85 0.55 0.60 0.15 0.60 0.65 0.60 0.60 0.45 0.40 0.70 0.30 0.65 0.35 0.60 0.55 0.60
40 0.40 0.50 0.20 0.40 0.30 0.40 0.30 0.45 0.55 0.50 0.25 0.30 0.70 0.65 0.30 0.25 0.60 0.25 0.55 0.40 0.25 0.65 0.70 0.80 0.35 0.45 0.25 0.55 0.35 0.30
41 0.20 0.60 0.40 0.50 0.45 0.60 0.35 0.25 0.55 0.20 0.45 0.50 0.85 0.45 0.50 0.45 0.65 0.45 0.20 0.30 0.35 0.55 0.80 0.20 0.35 0.50 0.65 0.65 0.80 0.40
42 0.20 0.40 0.20 0.35 0.15 0.05 0.30 0.30 0.55 0.35 0.05 0.20 0.70 0.45 0.20 0.40 0.65 0.60 0.25 0.10 0.05 0.30 0.50 0.65 0.35 0.60 0.50 0.60 0.60 0.40
43 0.50 0.50 0.50 0.35 0.40 0.30 0.35 0.35 0.15 0.35 0.25 0.45 0.20 0.60 0.50 0.35 0.35 0.05 0.30 0.30 0.15 0.30 0.60 0.40 0.55 0.40 0.20 0.55 0.70 0.40
44 0.90 0.90 0.95 1.00 1.00 0.85 0.80 0.90 0.95 0.75 0.85 0.85 0.95 0.80 1.00 1.00 0.95 0.85 0.85 0.65 0.95 0.85 0.95 0.90 0.95 0.95 0.95 0.95 1.00 0.95
45 0.55 0.40 0.60 0.80 0.75 0.60 0.45 0.55 0.60 0.70 0.70 0.55 0.45 0.40 0.30 0.55 0.15 0.55 0.40 0.65 0.30 0.65 0.50 0.35 0.55 0.65 0.45 0.60 0.75 0.60
46 0.00 0.00 0.00 0.05 0.35 0.10 0.00 0.15 0.00 0.10 0.10 0.05 0.05 0.10 0.00 0.00 0.00 0.10 0.10 0.00 0.05 0.05 0.05 0.20 0.00 0.05 0.10 0.10 0.00 0.10
47 0.45 0.35 0.25 0.25 0.20 0.20 0.10 0.20 0.15 0.60 0.25 0.45 0.45 0.30 0.30 0.15 0.30 0.30 0.15 0.35 0.20 0.40 0.40 0.30 0.20 0.25 0.10 0.55 0.45 0.30
48 0.00 0.15 0.30 0.10 0.00 0.15 0.20 0.05 0.20 0.10 0.10 0.25 0.05 0.10 0.25 0.15 0.00 0.25 0.20 0.10 0.20 0.05 0.10 0.10 0.15 0.20 0.10 0.10 0.05 0.20
49 0.05 0.05 0.25 0.25 0.30 0.15 0.10 0.20 0.25 0.25 0.40 0.30 0.15 0.25 0.05 0.05 0.00 0.15 0.25 0.30 0.00 0.00 0.00 0.15 0.15 0.05 0.25 0.15 0.05 0.10
50 0.25 0.05 0.25 0.00 0.05 0.10 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.05 0.10 0.05 0.05 0.00 0.00 0.05 0.00 0.05 0.05 0.05 0.10 0.00 0.00 0.05 0.15 0.05
51 0.05 0.15 0.05 0.10 0.15 0.05 0.15 0.15 0.20 0.15 0.20 0.15 0.15 0.10 0.05 0.05 0.20 0.05 0.35 0.15 0.35 0.15 0.20 0.30 0.15 0.00 0.10 0.05 0.10 0.10
52 0.00 0.00 0.00 0.00 0.00 0.05 0.05 0.10 0.00 0.00 0.05 0.05 0.10 0.00 0.15 0.10 0.00 0.00 0.05 0.05 0.00 0.05 0.15 0.10 0.00 0.00 0.10 0.15 0.00 0.05
53 0.05 0.05 0.10 0.15 0.00 0.05 0.20 0.20 0.05 0.00 0.10 0.20 0.05 0.25 0.05 0.05 0.00 0.00 0.25 0.15 0.25 0.00 0.00 0.00 0.05 0.15 0.10 0.10 0.05 0.15
54 0.05 0.20 0.05 0.05 0.05 0.00 0.05 0.05 0.05 0.25 0.05 0.05 0.05 0.05 0.10 0.20 0.10 0.00 0.20 0.20 0.05 0.05 0.00 0.00 0.00 0.00 0.00 0.05 0.15 0.05
55 0.10 0.05 0.05 0.00 0.00 0.00 0.10 0.05 0.00 0.05 0.00 0.00 0.00 0.00 0.05 0.00 0.00 0.00 0.05 0.05 0.05 0.00 0.05 0.05 0.00 0.00 0.00 0.00 0.10 0.00
56 0.05 0.15 0.10 0.10 0.10 0.05 0.15 0.15 0.30 0.20 0.10 0.25 0.10 0.10 0.05 0.10 0.20 0.00 0.25 0.15 0.10 0.05 0.00 0.05 0.20 0.15 0.10 0.10 0.15 0.20
57 0.10 0.00 0.05 0.10 0.10 0.05 0.05 0.20 0.00 0.15 0.10 0.05 0.10 0.15 0.00 0.05 0.05 0.10 0.20 0.25 0.10 0.00 0.00 0.20 0.10 0.00 0.20 0.20 0.05 0.05
58 0.05 0.10 0.20 0.00 0.05 0.10 0.10 0.25 0.10 0.25 0.20 0.15 0.20 0.15 0.15 0.05 0.05 0.15 0.05 0.15 0.05 0.15 0.10 0.10 0.05 0.10 0.20 0.15 0.05 0.10
59 0.05 0.10 0.10 0.05 0.10 0.10 0.05 0.15 0.00 0.05 0.10 0.05 0.05 0.10 0.25 0.05 0.15 0.25 0.05 0.10 0.05 0.10 0.00 0.10 0.15 0.15 0.00 0.00 0.20 0.05
60 0.55 0.65 0.70 0.30 0.60 0.45 0.50 0.50 0.40 0.40 0.35 0.65 0.20 0.75 0.55 0.35 0.35 0.50 0.40 0.40 0.50 0.45 0.50 0.75 0.20 0.50 0.55 0.45 0.65 0.90
61 0.05 0.15 0.05 0.00 0.10 0.00 0.00 0.10 0.05 0.05 0.10 0.10 0.05 0.10 0.05 0.00 0.00 0.10 0.15 0.05 0.20 0.10 0.15 0.20 0.05 0.15 0.05 0.15 0.05 0.15
62 0.45 0.40 0.40 0.45 0.20 0.20 0.25 0.30 0.25 0.55 0.15 0.70 0.40 0.50 0.45 0.30 0.10 0.50 0.40 0.15 0.05 0.20 0.15 0.35 0.25 0.35 0.50 0.40 0.10 0.35
63 0.10 0.00 0.05 0.15 0.05 0.00 0.15 0.20 0.15 0.05 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.10 0.00 0.00 0.00 0.00 0.00 0.00
64 0.10 0.05 0.05 0.00 0.05 0.00 0.00 0.05 0.05 0.00 0.05 0.15 0.00 0.00 0.00 0.00 0.00 0.00 0.05 0.05 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.10 0.05 0.00
65 0.05 0.00 0.15 0.20 0.05 0.10 0.00 0.05 0.05 0.10 0.05 0.05 0.05 0.00 0.05 0.00 0.05 0.05 0.05 0.10 0.10 0.10 0.00 0.10 0.15 0.05 0.05 0.05 0.05 0.05
66 0.20 0.35 0.05 0.05 0.30 0.15 0.40 0.40 0.50 0.20 0.15 0.25 0.10 0.20 0.05 0.25 0.10 0.25 0.15 0.20 0.30 0.20 0.20 0.20 0.10 0.10 0.20 0.35 0.40 0.25
67 0.00 0.00 0.05 0.05 0.00 0.15 0.05 0.05 0.05 0.05 0.00 0.05 0.05 0.00 0.00 0.10 0.05 0.00 0.00 0.05 0.00 0.00 0.10 0.00 0.05 0.10 0.10 0.00 0.00 0.00
68 0.05 0.40 0.20 0.30 0.25 0.15 0.30 0.40 0.60 0.25 0.20 0.25 0.20 0.25 0.15 0.20 0.25 0.15 0.40 0.50 0.40 0.35 0.30 0.35 0.15 0.15 0.35 0.30 0.15 0.30
69 0.15 0.00 0.20 0.10 0.25 0.05 0.00 0.00 0.00 0.10 0.05 0.10 0.30 0.05 0.50 0.20 0.20 0.00 0.10 0.20 0.10 0.30 0.15 0.25 0.30 0.30 0.10 0.25 0.35 0.10
70 0.20 0.45 0.05 0.20 0.20 0.20 0.30 0.35 0.20 0.20 0.30 0.25 0.20 0.20 0.30 0.20 0.30 0.25 0.20 0.20 0.20 0.25 0.55 0.40 0.25 0.15 0.50 0.40 0.20 0.40
71 0.35 0.25 0.10 0.20 0.20 0.35 0.35 0.20 0.15 0.15 0.30 0.05 0.15 0.15 0.30 0.25 0.10 0.30 0.35 0.05 0.45 0.35 0.10 0.15 0.05 0.10 0.25 0.20 0.45 0.30
72 0.10 0.35 0.25 0.05 0.10 0.40 0.05 0.05 0.25 0.05 0.20 0.10 0.20 0.35 0.35 0.10 0.10 0.30 0.05 0.20 0.05 0.25 0.25 0.50 0.15 0.40 0.30 0.05 0.30 0.25
73 0.20 0.05 0.25 0.10 0.05 0.15 0.05 0.00 0.10 0.20 0.15 0.15 0.00 0.15 0.00 0.25 0.10 0.25 0.15 0.05 0.15 0.05 0.05 0.25 0.20 0.25 0.05 0.10 0.10 0.30
74 0.30 0.40 0.25 0.20 0.25 0.40 0.20 0.30 0.15 0.40 0.15 0.25 0.35 0.30 0.25 0.15 0.10 0.25 0.45 0.35 0.15 0.40 0.25 0.45 0.40 0.50 0.20 0.50 0.15 0.25
75 0.20 0.30 0.15 0.20 0.10 0.25 0.10 0.20 0.10 0.05 0.20 0.15 0.20 0.20 0.20 0.25 0.15 0.20 0.10 0.10 0.10 0.20 0.35 0.55 0.20 0.20 0.15 0.10 0.20 0.25
76 0.15 0.15 0.40 0.75 0.50 0.20 0.40 0.40 0.35 0.45 0.25 0.30 0.00 0.30 0.05 0.00 0.00 0.00 0.60 0.20 0.30 0.05 0.25 0.40 0.10 0.05 0.10 0.10 0.05 0.05
77 0.35 0.35 0.10 0.05 0.05 0.30 0.25 0.50 0.35 0.15 0.00 0.40 0.15 0.15 0.50 0.30 0.00 0.10 0.15 0.20 0.20 0.15 0.25 0.30 0.55 0.45 0.55 0.35 0.15 0.40
78 0.10 0.20 0.25 0.00 0.20 0.15 0.10 0.00 0.00 0.10 0.10 0.05 0.10 0.40 0.20 0.05 0.00 0.10 0.05 0.05 0.00 0.20 0.05 0.05 0.20 0.30 0.10 0.15 0.00 0.25
79 0.15 0.40 0.35 0.25 0.20 0.40 0.15 0.45 0.40 0.35 0.35 0.20 0.35 0.25 0.35 0.35 0.00 0.20 0.30 0.15 0.30 0.25 0.15 0.60 0.50 0.65 0.25 0.35 0.05 0.25
80 0.15 0.30 0.40 0.20 0.40 0.45 0.25 0.20 0.35 0.10 0.05 0.15 0.40 0.20 0.20 0.20 0.00 0.15 0.30 0.25 0.50 0.25 0.40 0.45 0.60 0.20 0.30 0.40 0.00 0.30
81 0.15 0.10 0.15 0.00 0.05 0.20 0.25 0.15 0.05 0.10 0.05 0.10 0.05 0.20 0.05 0.15 0.00 0.05 0.00 0.00 0.05 0.25 0.25 0.25 0.00 0.15 0.15 0.20 0.25 0.20
82 0.20 0.35 0.45 0.55 0.30 0.45 0.15 0.45 0.25 0.40 0.60 0.65 0.70 0.45 0.60 0.20 0.00 0.15 0.35 0.45 0.50 0.25 0.45 0.65 0.45 0.65 0.15 0.40 0.30 0.45
83 0.10 0.35 0.40 0.10 0.15 0.20 0.10 0.15 0.10 0.30 0.00 0.10 0.20 0.15 0.20 0.05 0.10 0.15 0.40 0.20 0.15 0.15 0.00 0.25 0.15 0.25 0.20 0.15 0.10 0.20
84 0.05 0.20 0.20 0.15 0.25 0.05 0.20 0.15 0.40 0.35 0.20 0.10 0.30 0.20 0.05 0.05 0.05 0.10 0.50 0.25 0.20 0.20 0.05 0.20 0.20 0.05 0.20 0.05 0.15 0.00
85 0.05 0.00 0.05 0.00 0.10 0.05 0.05 0.20 0.00 0.20 0.10 0.15 0.00 0.05 0.05 0.00 0.00 0.00 0.20 0.10 0.05 0.00 0.00 0.05 0.10 0.05 0.00 0.05 0.05 0.00
86 0.05 0.25 0.15 0.20 0.60 0.25 0.00 0.05 0.00 0.15 0.25 0.30 0.25 0.25 0.30 0.20 0.05 0.15 0.15 0.20 0.10 0.10 0.00 0.45 0.05 0.20 0.05 0.40 0.30 0.30
87 0.40 0.65 0.65 0.25 0.55 0.40 0.30 0.20 0.20 0.40 0.30 0.35 0.55 0.65 0.55 0.25 0.50 0.40 0.45 0.30 0.30 0.30 0.45 0.90 0.35 0.65 0.70 0.50 0.45 0.55
88 0.40 0.20 0.55 0.20 0.55 0.35 0.45 0.65 0.65 0.55 0.60 0.55 0.45 0.50 0.50 0.40 0.20 0.50 0.55 0.25 0.20 0.75 0.45 0.40 0.35 0.65 0.30 0.45 0.55 0.50
89 0.10 0.00 0.05 0.05 0.10 0.10 0.05 0.20 0.05 0.20 0.05 0.05 0.00 0.10 0.00 0.10 0.25 0.05 0.10 0.15 0.05 0.00 0.25 0.30 0.30 0.05 0.10 0.20 0.20 0.10